### v1

# Sequence Reconstruction for Single-Deletion Single-Substitution Channel

Abstract
The central problem in sequence reconstruction is to find the minimum number of distinct channel outputs required to uniquely reconstruct the transmitted sequence. According to Levenshtein’s work in 2001, this number is determined by the maximum size of the intersection between the error balls of any two distinct input sequences of the channel. In this work, we study the sequence reconstruction problem for single-deletion single-substitution channel, assuming that the transmitted sequence belongs to a $q$-ary code with minimum Hamming distance at least $2$, where $q\geq 2$ is any fixed integer. Specifically, we prove that for any two $q$-ary sequences with Hamming distance $d\geq 2$, the size of the intersection of their error balls is upper bounded by $2qn-3q-2-\delta_{q,2}$, where $\delta_{i,j}$ is the Kronecker delta. We also prove the tightness of this bound by constructing two sequences the intersection size of whose error balls achieves this bound.
Index Terms:
Sequence reconstruction, reconstruction codes, deletion, substitution.
I Introduction
We consider a communication scenario where a codeword $\bm{x}$ from some codebook $\mathscr{C}$ is transmitted over a number of identical channels and the goal is to reconstruct $\bm{x}$ from all (erroneous) channel outputs (also referred to as reads in data storage applications). This problem, also known as the sequence reconstruction problem, was first proposed by Levenshtein [1], [2], and in recent years, gained more and more attentions due to its applications in DNA data storage [3]. The central problem in sequence reconstruction is to determine the minimum number of distinct channel outputs (reads) required to uniquely reconstruct $\bm{x}$. This number was shown to be equal to one plus the maximum size of the intersection between the error balls of any two distinct codewords of $\mathscr{C}$ (also referred to as the read coverage of $\mathscr{C}$ for the channel) [1]. Therefore, deriving the read coverage of $\mathscr{C}$ is essential for solving the sequence reconstruction problem. On the other hand, designing of codes with given read coverage, called reconstruction codes, is also an interesting problem for sequence reconstruction.
In his seminal work [1], Levenshtein studied the sequence reconstruction problem for deletion, insertion, substitution and transposition separately, where $\mathscr{C}$ is taken to be the set of all $q$-ary sequence. For the more general case that $\mathscr{C}$ is an $(\ell-1)$-deletion correcting code for some positive integer $\ell\leq t$, the problem was studied in [4] and [5] for $t$-deletion channel, and in [6] for $t$-insertion channel. Reconstruction codes for two-deletion channels can be found in [7, 8] and reconstruction codes for two-insertion channels can be found in [9]. Reconstruction codes for $q$-ary single-edit channel $(q\geq 2)$ was constructed in [10] by generalizing the construction in [11], where an edit error means a deletion, an insertion or a substitution error. Reconstruction codes for single-burst-insertion/deletion were constructed in [12], where a burst of $t$ deletions/insertions means $t$ deletions or $t$ insertions occurring at consecutive positions. In these constructions, each reads allows only one type of error.
In practical applications, a read may suffers from different error types, for example, both a deletion and an insertion, or both a deletion and a substitution. It was shown in [13] that a code $\mathscr{C}$ can correct $t$ deletions if and only if it can correct $t$ insertions. However, the intersection size of $t$-deletion balls of two sequences is not necessarily equal to the intersection size of their $t$-insertion balls when the intersections are not empty. Therefore, unlike the classic error correction problem, in sequence reconstruction problem, the deletion channel and the insertion channel must be treated separately. The reconstruction problem for single-insertion single-substitution was studied in [14], where the maximum intersection size of single-insertion single-substitution balls was proved to be $\lfloor\frac{n-2}{2}\rfloor\lceil\frac{n-2}{2}\rceil+4n$. The size of single-deletion multiple-substitution ball was also computed in [14], but their intersection size was not considered. In a more recent work [15], the size of the error ball for channels with multiple types of errors and at most three edits was studied. To the best of our knowledge, deriving the maximum intersection size of single-deletion single-substitution balls is still an open problem.
In this work, we study the sequence reconstruction problem for $q$-ary single-deletion single-substitution channel, where $q\geq 2$ is an arbitrarily fixed integer. ((((For example, in DNA data storage, we take $q=4$.)))) We prove that for any two $q$-ary sequences with Hamming distance $d\geq 2$, the size of the intersection of their error balls is upper bounded by $2qn-3q-2-\delta_{q,2}$, where $\delta_{i,j}$ is the Kronecker delta. We also show that there exist two sequences the intersection size of whose error balls achieves this bound, which proves that the bound is tight. Note that the requirement that the Hamming distance between two sequences is at least $2$ can be satisfied by adding one parity check symbol, so results in one symbol of redundancy.
The single-deletion single-substitution channel was studied in several existing works under error correction model or list-decoding model [16]--[21]. The best known single-deletion single-substitution correcting binary codes has $4\log n+o(\log n)$ bits of redundancy, where $n$ is the code length [21]. By our result, when the number of reads is $2qn-3q-1-\delta_{q,2}$, one symbol of redundancy is sufficient to guarantee correct reconstruction of the transmitted sequence.
The paper is organized as follows. In Section II, we describe the problem and our main result, as well as some simple discussions that will help to prove our main result. In Section III, we develop a method for finding the intersection size of two error balls. We give a formal proof of our main result in Section IV and make conclusions and some discussions for future work in Section V.
II Problem description and main result
For any integers $m\leq n$, let $[m,n]=\{m,m+1,\ldots,n\}$ and called it an interval. If $m=1$, let $[n]=[1,n]$. If $A$ is a finite set of integers, then $|A|$ is the size (cardinality) of $A$, and $\min(A)~{}($resp. $\max(A))$ is the smallest ((((resp. greatest)))) number in $A$. For simplicity, we denote $A\backslash i=A\backslash\{i\}$ for any $i\in A$. If we denote $A=\{i_{1},i_{2},\cdots,i_{k}\}$, we always assume that $i_{1}<i_{2}<\cdots<i_{k}$.
Let $\Sigma_{q}=\{0,1,\cdots,q-1\}$, where $q\geq 2$ is an arbitrarily fixed integer. For any $\bm{x}\in\Sigma_{q}^{n}$, let $x_{i}$ denote the $i$th component of $\bm{x}$ and denote $\bm{x}=x_{1}x_{2}\cdots x_{n}$ or $\bm{x}=(x_{1},x_{2},\cdots,x_{n})$. If $D=\{i_{1},i_{2},\cdots,i_{m}\}\subseteq[n]$, let $x_{D}=x_{i_{1}}x_{i_{2}}\cdots x_{i_{m}}$ and call it a subsequence of $\bm{x}$. If $D$ is an interval, $x_{D}$ is called a substring of $\bm{x}$. A run of $\bm{x}$ is a maximal substring of $\bm{x}$ consisting of identical symbols. For any two given distinct symbols $a,b\in\Sigma_{q}$, let $A_{n}(ab)$ denote the alternating sequence of length $n$ that starts with $a$ and consists of $a,b$. For example, $A_{5}(01)=01010$ and $A_{6}(01)=010101$.
For any $\bm{x},\bm{x}^{\prime}\in\Sigma_{q}^{n}$, the Hamming distance between $\bm{x}$ and $\bm{x}^{\prime}$, denoted by $d_{\text{H}}(\bm{x},\bm{x}^{\prime})$, is defined as the number of $i\in[n]$ such that $x_{i}\neq x^{\prime}_{i}$. The Levenshtein distance between $\bm{x}$ and $\bm{x}^{\prime}$, denoted by $d_{\text{L}}(\bm{x},\bm{x}^{\prime})$, is defined as the smallest integer $\ell$ such that $\bm{x},\bm{x}^{\prime}$ share some subsequence of length $n-\ell$.
Let $t$ and $s$ be non-negative integers such that $t+s<n$. For any $\bm{x}\in\Sigma_{q}^{n}$, the $t$-deletion $s$-substitution ball of $\bm{x}$, denoted by $B^{\text{D,S}}_{t,s}(\bm{x})$, is the set of all sequences that can be obtained from $\bm{x}$ by exact $t$ deletions and at most $s$ substitutions. The $t$-deletion ball of $\bm{x}$ is $B^{\text{D}}_{t}(\bm{x})\triangleq B^{\text{D,S}}_{t,0}(\bm{x})$, and the $s$-substitution ball of $\bm{x}$ is $B^{\text{S}}_{s}(\bm{x})\triangleq B^{\text{D,S}}_{0,s}(\bm{x})$. For $B\in\{B^{\text{D}}_{t},B^{\text{S}}_{s},B^{\text{D,S}}_{t,s}\}$, let
$B(\bm{x},\bm{x}^{\prime})\triangleq B(\bm{x})\cap B(\bm{x}^{\prime}).$
Given a code $\mathscr{C}\subseteq\Sigma_{q}^{n}$, let
$\nu(\mathscr{C};B)\triangleq\max\{|B(\bm{x},\bm{x}^{\prime})|:\bm{x},\bm{x}^{\prime}\in\mathscr{C},\bm{x}\neq\bm{x}^{\prime}\}$
called the read coverage of $\mathscr{C}$. A central problem in sequence reconstruction is to compute $\nu(\mathscr{C};B)$ for the given $\mathscr{C}$ and $B$. Another problem is, for the given error ball function $B$ and positive integer $N$, to design a code $\mathscr{C}\subseteq\Sigma_{q}^{n}$ with $\nu(\mathscr{C};B)<N$, called an $(n,N,B)$-reconstruction code.
In this work, we assume $q\geq 2$ is a fixed positive integer. We consider the sequence reconstruction problem for $q$-ary single-deletion single-substitution channel ((((i.e., $B=B^{\text{D,S}}_{1,1})$. Our main result is the following theorem.
Theorem 1
Suppose $n\geq\max\{\frac{q+23}{2},\frac{5q+19}{q-1}\}$. For any $\bm{x},\bm{x}^{\prime}\in\Sigma_{q}^{n}$ with $d_{\text{H}}(\bm{x},\bm{x}^{\prime})\geq 2$, we have
$|B^{\text{D,S}}_{1,1}(\bm{x},\bm{x}^{\prime})|\leq 2qn-3q-2-\delta_{q,2}$
where $\delta_{i,j}$ is the Kronecker delta. Moreover, there exist two sequences $\bm{x},\bm{x}^{\prime}\in\Sigma_{q}^{n}$ with $d_{\text{H}}(\bm{x},\bm{x}^{\prime})=2$ and $|B^{\text{D,S}}_{1,1}(\bm{x},\bm{x}^{\prime})|=2qn-3q-2-\delta_{q,2}$.
The proof of Theorem 1 will be given in Section IV. In the rest of this section, we state some simple results that will be used by the proof.
II-A Intersection size of error balls of $q$-ary substitution channel
First, the size of the $q$-ary substitution ball satisfies ((((e.g., see [22, Chapter 1]))))
$|B^{\text{S}}_{s}(\bm{x})|=\sum_{k=0}^{s}\binom{n}{k}(q-1)^{k},~{}\forall\!~{}\bm{x}\in\Sigma_{q}^{n}.$
In particular, for $s=1$, we have
$|B^{\text{S}}_{1}(\bm{x})|=1+(q-1)n,~{}\forall\!~{}\bm{x}\in\Sigma_{q}^{n}.$ (1)
For the intersection size of single-substitution balls, we have the following simple remark.
Remark 1
For any $\bm{x},\bm{x}^{\prime}\in\Sigma_{q}^{n}$, we have
$|B^{\text{S}}_{1}(\bm{x},\bm{x}^{\prime})|=\!\left\{\!\begin{aligned} &q,~{}\text{if}~{}d_{\text{H}}(\bm{x},\bm{x}^{\prime})=1;\\ &2,~{}\text{if}~{}d_{\text{H}}(\bm{x},\bm{x}^{\prime})=2;\\ &0,~{}\text{if}~{}d_{\text{H}}(\bm{x},\bm{x}^{\prime})\geq 3.\end{aligned}\right.$
II-B Some useful observations and lemma
Consider the intersection size of $t$-deletion $s$-substitution balls. Suppose $\bm{x},\bm{x}^{\prime}\in\Sigma_{q}^{n}$. By the definition of $B^{\text{D,S}}_{t,s}$, it is easy to see that
$B^{\text{D,S}}_{t,s}(\bm{x},\bm{x}^{\prime})=\bigcup_{\bm{z}\in B^{\text{D}}_{t}(\bm{x}),\bm{z}^{\prime}\in B^{\text{D}}_{t}(\bm{x}^{\prime})}B^{\text{S}}_{s}(\bm{z},\bm{z}^{\prime}).$
For the special case that $t=s=1$, we have
$B^{\text{D,S}}_{1,1}(\bm{x},\bm{x}^{\prime})$ $=\bigcup_{\bm{z}\in B^{\text{D}}_{1}(\bm{x}),\bm{z}^{\prime}\in B^{\text{D}}_{1}(\bm{x}^{\prime})}B^{\text{S}}_{1}(\bm{z},\bm{z}^{\prime})$
$=\bigcup_{j,j^{\prime}\in[n]}B^{\text{S}}_{1}(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}}).$
Note that by Remark 1, $|B^{\text{S}}_{1}(\bm{z},\bm{z}^{\prime})|=0$ when $d_{\text{H}}(\bm{z},\bm{z}^{\prime})\geq 3$. Then we have the following observation:
Observation 1: It holds that
$B^{\text{D,S}}_{1,1}(\bm{x},\bm{x}^{\prime})=\bigcup_{(\bm{z},\bm{z}^{\prime})\in\Lambda}B^{\text{S}}_{1}(\bm{z},\bm{z}^{\prime})$
where
$\Lambda=\Lambda(\bm{x},\bm{x}^{\prime})$$\triangleq\Big{\{}(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}}):j,j^{\prime}\in[n]~{}\text{and}~{}$
$~{}~{}~{}~{}~{}~{}d_{\text{H}}(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}})\leq 2\Big{\}}.$ (2)
To compute $d_{\text{H}}(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}})$, we introduce some notations as follows. Let
$S=S(\bm{x},\bm{x}^{\prime})\triangleq\{i\in[n]:x_{i}\neq x^{\prime}_{i}\}.$ (3)
Then we have $|S|=d_{\text{H}}(\bm{x},\bm{x}^{\prime})$, and so we can denote
$S=\{i_{1},i_{2},\cdots,i_{d}\}$
where $d=d_{\text{H}}(\bm{x},\bm{x}^{\prime})$ and $i_{1}<i_{2}<\cdots<i_{d}$ according to our previous convention. Moreover, we let
$T^{L}=T^{L}(\bm{x},\bm{x}^{\prime})\triangleq\{i\in[2,n]:x_{i}\neq x^{\prime}_{i-1}\}.$ (4)
and
$T^{R}=T^{R}(\bm{x},\bm{x}^{\prime})\triangleq\{i\in[2,n]:x_{i-1}\neq x^{\prime}_{i}\}.$ (5)
From these definitions, it is easy to see that
$T^{R}(\bm{x},\bm{x}^{\prime})=T^{L}(\bm{x}^{\prime},\bm{x})$
We remind the reader that in the notations $T^{L}(\bm{x},\bm{x}^{\prime})$ and $T^{R}(\bm{x},\bm{x}^{\prime})$, $(\bm{x},\bm{x}^{\prime})$ is viewed as an ordered pair. Note that by the definitions, $T^{L}=T^{L}(\bm{x},\bm{x}^{\prime})\neq T^{L}(\bm{x}^{\prime},\bm{x})$ and $T^{R}=T^{R}(\bm{x},\bm{x}^{\prime})\neq T^{R}(\bm{x}^{\prime},\bm{x})$.
Now, we have the second useful observation.
Observation 2: For any $j,j^{\prime}\in[n]$, $j\leq j^{\prime}$, we have
$d_{\text{H}}(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}})$
$=\left|\big{(}S\cap[1,j-1]\big{)}\cup\big{(}T^{L}\cap[j+1,j^{\prime}]\big{)}\cup\big{(}S\cap[j^{\prime}+1,n]\big{)}\right|$
$=|S\cap[1,j-1]|+|T^{L}\cap[j+1,j^{\prime}]|+|S\cap[j^{\prime}+1,n]|;$
and
$d_{\text{H}}(x_{[n]\backslash j^{\prime}},x^{\prime}_{[n]\backslash j})$
$=\left|\big{(}S\cap[1,j-1]\big{)}\cup\big{(}T^{R}\cap[j+1,j^{\prime}]\big{)}\cup\big{(}S\cap[j^{\prime}+1,n]\big{)}\right|$
$=|S\cap[1,j-1]|+|T^{R}\cap[j+1,j^{\prime}]|+|S\cap[j^{\prime}+1,n]|.$
The following lemma will be used to exclude repeat count of sequence pairs in $\Lambda$.
Lemma 1
Suppose $j_{1},j_{2},j_{1}^{\prime},j_{2}^{\prime}\in[n]$ such that $j_{1}\leq j_{2}$ and $j_{1}^{\prime}\leq j_{2}^{\prime}$. The following hold.
  • 1)
    If $[j_{1},j_{2}-1]\cap S=[j_{1}+1,j_{2}]\cap T^{L}=\emptyset$, then $x_{[j_{1},j_{2}]}$ is contained in a run of $\bm{x}$.
  • 2)
    If $[j^{\prime}_{1}+1,j^{\prime}_{2}]\cap S=[j^{\prime}_{1}+1,j^{\prime}_{2}]\cap T^{L}=\emptyset$, then $x^{\prime}_{[j^{\prime}_{1},j^{\prime}_{2}]}$ is contained in a run of $\bm{x}^{\prime}$.
proof 1
We first prove 1). If $[j_{1},j_{2}-1]\cap S=\emptyset$, then by the definition of $S$, we have $x_{i}=x^{\prime}_{i}$ for all $i\in[j_{1},j_{2}-1]$; if $[j_{1}+1,j_{2}]\cap T^{L}=\emptyset$, then by the definition of $T^{L}$, we have $x_{i}=x^{\prime}_{i-1}$ for all $i\in[j_{1}+1,j_{2}]$. Hence, we can obtain $x_{i}=x_{i}^{\prime}=x_{i+1}$ for all $i\in[j_{1},j_{2}-1]$, which implies that $x_{[j_{1},j_{2}]}$ is contained in a run of $\bm{x}$.
The proof of 2) is similar to 1). From the assumption that $[j^{\prime}_{1}+1,j^{\prime}_{2}]\cap S=[j^{\prime}_{1}+1,j^{\prime}_{2}]\cap T^{L}=\emptyset$, we can obtain $x^{\prime}_{i}=x_{i}=x^{\prime}_{i-1}$ for all $i\in[j^{\prime}_{1}+1,j^{\prime}_{2}]$, which implies that $x^{\prime}_{[j^{\prime}_{1},j^{\prime}_{2}]}$ is contained in a run of $\bm{x}^{\prime}$.
By Lemma 1, if two pairs $(j_{1},j_{2})$ and $(j^{\prime}_{1},j^{\prime}_{2})$ such that $j_{1},j_{2},j_{1}^{\prime},j_{2}^{\prime}$ satisfy the conditions of Lemma 1, then for any $(j,j^{\prime})\in[j_{1},j_{2}]\times[j_{1}^{\prime},j_{2}^{\prime}]$, we have $(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}})=(x_{[n]\backslash j_{1}},x^{\prime}_{[n]\backslash j_{1}^{\prime}})=(x_{[n]\backslash j_{2}},x^{\prime}_{[n]\backslash j_{2}^{\prime}})$.
II-C The notation $\phi$
For $a\in\Sigma_{q}$ and $j_{1}\neq j_{2}\in[n]$, let $\phi^{j_{1}}_{j_{2};a}(\bm{x})$ be the sequence obtained from $\bm{x}$ by deleting $x_{j_{1}}$ and substituting $x_{j_{2}}$ with $a$. For example, if $\bm{x}=10212201$, then $\phi^{6}_{3;0}(\bm{x})=1001201$.
In our subsequent discussions, it will be helpful to describe $B^{\text{S}}_{1}(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}})$ using the notation $\phi$.
Example 1
Suppose $\bm{x},\bm{x}^{\prime}\in\Sigma_{2}^{8}$ such that
$\bm{x}$ $=01010111$
$~{}\bm{x}^{\prime}$ $=01101011.$
Then we can find
$x_{[n]\backslash 4}$ $=0100111$
$~{}\bm{x}^{\prime}_{[n]\backslash 7}$ $=0110101$
and
$B^{\text{S}}_{1}(x_{[n]\backslash 4},x^{\prime}_{[n]\backslash 7})=\{0100101,0110111\}.$
In fact, we can easily check that $\phi^{4}_{7;x^{\prime}_{6}}(\bm{x})=0100101=\phi^{7}_{3;x_{3}}(\bm{x}^{\prime})$ and $\phi^{4}_{3;x^{\prime}_{3}}(\bm{x})=0110111=\phi^{7}_{6;x_{7}}(\bm{x}^{\prime})$.
In general, we have the following two remarks.
Remark 2
For any $j,j^{\prime}\in[n]$ such that $j\leq j^{\prime}$, if $d_{\text{H}}(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}})=2$, then by Observation 2, we can let $\big{(}S\cap[1,j-1]\big{)}\cup\big{(}T^{L}\cap[j+1,j^{\prime}]\big{)}\cup\big{(}S\cap[j^{\prime}+1,n]\big{)}=\{j_{1},j_{2}\}$. For each $\ell\in\{1,2\}$: if $j_{\ell}\in[1,j-1]\cup[j^{\prime}+1,n]$, let $\bm{z}_{\ell}=\phi^{j}_{j_{\ell};x^{\prime}_{j_{\ell}}}(\bm{x})$ and $\bm{w}_{\ell}=\phi^{j^{\prime}}_{j_{\ell};x_{j_{\ell}}}(\bm{x}^{\prime})$; if $j_{\ell}\in[j+1,j^{\prime}]$, let $\bm{z}_{\ell}=\phi^{j}_{j_{\ell};x^{\prime}_{j_{\ell}-1}}(\bm{x})$ and $\bm{w}_{\ell}=\phi^{j^{\prime}}_{j_{\ell}-1;x_{j_{\ell}}}(\bm{x}^{\prime})$. Then we have
$B^{\text{S}}_{1}(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}})=\{\bm{z}_{1},\bm{z}_{2}\}=\{\bm{w}_{1},\bm{w}_{2}\}.$
Similar results can be obtained when $d_{\text{H}}(x_{[n]\backslash j^{\prime}},x^{\prime}_{[n]\backslash j})=2$.
Remark 3
Similar to Remark 2, for any $j,j^{\prime}\in[n]$ such that $j\leq j^{\prime}$ and $d_{\text{H}}(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}})=1$, then by Observation 2, we can let $\big{(}S\cap[1,j-1]\big{)}\cup\big{(}T^{L}\cap[j+1,j^{\prime}]\big{)}\cup\big{(}S\cap[j^{\prime}+1,n]\big{)}=\{j_{1}\}$. If $j_{1}\in[1,j-1]\cup[j^{\prime}+1,n]$, then we have
$B^{\text{S}}_{1}(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}})$ =$=\big{\{}\phi^{j}_{j_{1};a}(\bm{x}):a\in\Sigma_{q}\big{\}}$
=$=\big{\{}\phi^{j^{\prime}}_{j_{1};a}(\bm{x}^{\prime}):a\in\Sigma_{q}\big{\}};$
if $j_{1}\in[j+1,j^{\prime}]$, then we have
$B^{\text{S}}_{1}(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}})$ =$=\big{\{}\phi^{j}_{j_{1};a}(\bm{x}):a\in\Sigma_{q}\big{\}}$
=$=\big{\{}\phi^{j^{\prime}}_{j_{1}-1;a}(\bm{x}^{\prime}):a\in\Sigma_{q}\big{\}}.$
Similar results can be obtained when $d_{\text{H}}(x_{[n]\backslash j^{\prime}},x^{\prime}_{[n]\backslash j})=1$.
III Methodology
In this section, we will always assume that $\bm{x},\bm{x}^{\prime}\in\Sigma_{q}^{n}$ are arbitrarily chosen such that $d=d_{\text{H}}(\bm{x},\bm{x}^{\prime})\geq 2$. By Observation 1, we have $B^{\text{D,S}}_{1,1}(\bm{x},\bm{x}^{\prime})=\bigcup_{(\bm{z},\bm{z}^{\prime})\in\Lambda}B^{\text{S}}_{1}(\bm{z},\bm{z}^{\prime})$, where $\Lambda=\Lambda(\bm{x},\bm{x}^{\prime})$ is defined by (II-B). To find the size of $B^{\text{D,S}}_{1,1}(\bm{x},\bm{x}^{\prime})$, we will develop a method to divide the set $\Lambda$, and correspondingly the set $B^{\text{D,S}}_{1,1}(\bm{x},\bm{x}^{\prime})$, into some subsets such that the size of each subset can be easily obtained.
Specifically, define some subsets as follows:
  • For each $\ell\in\{0,1,2\}$, let
    $\Lambda_{\ell}=\Lambda_{\ell}(\bm{x},\bm{x}^{\prime})$$\triangleq\left\{(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}}):j,j^{\prime}\in[n]\right.$
    $~{}~{}~{}~{}~{}~{}\left.\text{and}~{}d_{\text{H}}(x_{[n]\backslash j^{\prime}},x^{\prime}_{[n]\backslash j})=\ell\right\}$
    and
    $\Omega_{\ell}=\Omega_{\ell}(\bm{x},\bm{x}^{\prime})\triangleq\bigcup_{(\bm{z},\bm{z}^{\prime})\in\Lambda_{\ell}}B^{\text{S}}_{1}(\bm{z},\bm{z}^{\prime}).$
  • For each $\ell\in\{0,1,2\}$, let
    $\Lambda^{L}_{\ell}=\Lambda^{L}_{\ell}(\bm{x},\bm{x}^{\prime})$$\triangleq\left\{(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}}):(j,j^{\prime})\in[n]\times[n],\right.$
    $~{}~{}~{}~{}~{}\left.j\leq j^{\prime}~{}\text{and}~{}d_{\text{H}}(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}})=\ell\right\}$
    and
    $\Omega^{L}_{\ell}=\Omega^{L}_{\ell}(\bm{x},\bm{x}^{\prime})\triangleq\bigcup_{(\bm{z},\bm{z}^{\prime})\in\Lambda^{L}_{\ell}}B^{\text{S}}_{1}(\bm{z},\bm{z}^{\prime}).$
  • For each $\ell\in\{0,1,2\}$, let
    $\Lambda^{R}_{\ell}=\Lambda^{R}_{\ell}(\bm{x},\bm{x}^{\prime})$$\triangleq\left\{(x_{[n]\backslash j^{\prime}},x^{\prime}_{[n]\backslash j}):(j,j^{\prime})\in[n]\times[n],\right.$
    $~{}~{}~{}~{}~{}\left.j\leq j^{\prime}~{}\text{and}~{}d_{\text{H}}(x_{[n]\backslash j^{\prime}},x^{\prime}_{[n]\backslash j})=\ell\right\}$
    and
    $\Omega^{R}_{\ell}=\Omega^{R}_{\ell}(\bm{x},\bm{x}^{\prime})\triangleq\bigcup_{(\bm{z},\bm{z}^{\prime})\in\Lambda^{R}_{\ell}}B^{\text{S}}_{1}(\bm{z},\bm{z}^{\prime}).$
By the above definitions, we have
$\Lambda_{\ell}=\Lambda^{L}_{\ell}\cup\Lambda^{R}_{\ell}$
for each $\ell\in\{0,1,2\}$ and
$\Lambda=\bigcup_{\ell=0}^{2}\Lambda_{\ell}=\bigcup_{\ell=0}^{2}(\Lambda^{L}_{\ell}\cup\Lambda^{R}_{\ell}).$
Correspondingly, we have
$\Omega_{\ell}=\Omega^{L}_{\ell}\cup\Omega^{R}_{\ell}$
for each $\ell\in\{0,1,2\}$ and
$B^{\text{D,S}}_{1,1}(\bm{x},\bm{x}^{\prime})=\bigcup_{\ell=0}^{2}\Omega_{\ell}=\bigcup_{\ell=0}^{2}(\Omega^{L}_{\ell}\cup\Omega^{R}_{\ell}).$
Note that $\Lambda^{L}_{\ell}$ and $\Lambda^{R}_{\ell}$ are not necessarily disjoint, and so $\Omega^{L}_{\ell}$ and $\Omega^{R}_{\ell}$ are not necessarily disjoint.
We remark that $(\bm{x},\bm{x}^{\prime})$ should be viewed as an ordered pair in the notations $\Lambda^{X}_{\ell}(\bm{x},\bm{x}^{\prime})$, where $X\in\{L,R\}$ and $\ell\in\{0,1,2\}$. Moreover, by the definitions, $(\bm{z},\bm{z}^{\prime})\in\Lambda^{R}_{\ell}=\Lambda^{R}_{\ell}(\bm{x},\bm{x}^{\prime})$ if and only if $(\bm{z}^{\prime},\bm{z})\in\Lambda^{L}_{\ell}(\bm{x}^{\prime},\bm{x})$.
In the following three subsections, we will determine the set $\Lambda^{X}_{\ell}$ for each $X\in\{L,R\}$ and each $\ell\in\{0,1,2\}$.
III-A For $\Lambda^{L}_{0}(\bm{x},\bm{x}^{\prime})$ and $\Lambda^{R}_{0}(\bm{x},\bm{x}^{\prime})$
We first consider $\Lambda^{L}_{0}(\bm{x},\bm{x}^{\prime})$. Let $S$ and $T^{L}$ be defined according to (3) and (4), respectively.
Claim $\bm{0}$: Suppose $d=d_{\text{H}}(\bm{x},\bm{x}^{\prime})\geq 2$.
  • 1)
    If $|T^{L}\cap[i_{1}+1,i_{d}]|=0$, then $x_{[n]\backslash i_{1}}=x^{\prime}_{[n]\backslash i_{d}}$ and $\Lambda^{L}_{0}=\Lambda^{L}_{0}(\bm{x},\bm{x}^{\prime})=\{(x_{[n]\backslash i_{1}},x^{\prime}_{[n]\backslash i_{d}})\}$. Hence, we have $\Omega^{L}_{0}=\Omega^{L}_{0}(\bm{x},\bm{x}^{\prime})=B^{\text{S}}_{1}(x_{[n]\backslash i_{1}})=B^{\text{S}}_{1}(x^{\prime}_{[n]\backslash i_{d}})$.
  • 2)
    If $|T^{L}\cap[i_{1}+1,i_{d}]|\geq 1$, then $\Lambda^{L}_{0}(\bm{x},\bm{x}^{\prime})=\emptyset$.
proof 2
Let $k_{a}=\max(T^{L}\cap[1,i_{1}])$ if $T^{L}\cap[1,i_{1}]\neq\emptyset$, and $k_{a}=1$ otherwise. Similarly, let $k^{\prime}_{a}=\min(T^{L}\cap[i_{d}+1,n])-1$ if $T^{L}\cap[i_{d}+1,n]\neq\emptyset$, and $k^{\prime}_{a}=n$ otherwise. Then
$k_{a}\leq i_{1}<i_{d}\leq k^{\prime}_{a}.$
By the definition, to find $\Lambda^{L}_{0}(\bm{x},\bm{x}^{\prime})$, we need to find all $(j,j^{\prime})\in[n]\times[n]$ such that $j\leq j^{\prime}$ and $d_{\text{H}}(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}})=0$. By Observation 2, $d_{\text{H}}(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}})=0$ if and only if
$|S\cap[1,j-1]|+|T^{L}\cap[j+1,j^{\prime}]|+|S\cap[j^{\prime}+1,n]|=0,$
or equivalently,
$|S\cap[1,j-1]|=|T^{L}\cap[j+1,j^{\prime}]|=|S\cap[j^{\prime}+1,n]|=0.$ (6)
Note that $j\leq j^{\prime}$ and $S=\{i_{1},\cdots,i_{d}\}$ such that $i_{1}<\cdots<i_{d}$. Then from the conditions $|S\cap[1,j-1]|=|S\cap[j^{\prime}+1,n]|=0$, we have $j\leq i_{1}$ and $j^{\prime}\geq i_{d}$, which implies $T^{L}\cap[i_{1}+1,i_{d}]\subseteq T^{L}\cap[j+1,j^{\prime}]$. Further, from the condition $|T^{L}\cap[j+1,j^{\prime}]|=0$, we have $|T^{L}\cap[i_{1}+1,i_{d}]|\leq|T^{L}\cap[j+1,j^{\prime}]|=0$. Thus, if $|T^{L}\cap[i_{1}+1,i_{d}]|\geq 1$, then there is no $(j,j^{\prime})$ that satisfies the conditions in (6), and so $\Lambda^{L}_{0}(\bm{x},\bm{x}^{\prime})=\emptyset$.
Conversely, if $|T^{L}\cap[i_{1}+1,i_{d}]|=0$, then by the definition of $k_{a},k^{\prime}_{a},T^{L}$ and $S$, it is not hard to see that $x_{[n]\backslash i_{1}}=x^{\prime}_{[n]\backslash i_{d}}$ and $(j,j^{\prime})$ satisfies (6) if and only if $k_{a}\leq j\leq i_{1}<i_{d}\leq j^{\prime}\leq k^{\prime}_{a}$. Therefore, we have $\Lambda^{L}_{0}(\bm{x},\bm{x}^{\prime})=\{(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}}):k_{a}\leq j\leq i_{1}<i_{d}\leq j^{\prime}\leq k^{\prime}_{a}\}\neq\emptyset.$ Moreover, by the definition of $k_{a}$ and $k^{\prime}_{a}$, we can obtain $[k_{a},i_{1}-1]\cap S=\emptyset$ and $[k_{a}+1,i_{1}]\cap T^{L}=\emptyset$, so by 1) of Lemma 1, $x_{[k_{a},i_{1}]}$ is contained in a run of $\bm{x}$. Similarly, we can obtain $[i_{d}+1,k^{\prime}_{a}]\cap S=[i_{d}+1,k^{\prime}_{a}]\cap T^{L}=\emptyset$, and so by 2) of Lemma 1, $x^{\prime}_{[i_{d},k^{\prime}_{a}]}$ is contained in a run of $\bm{x}^{\prime}$. Thus, we have $\Lambda^{L}_{0}(\bm{x},\bm{x}^{\prime})=\{(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}}):k_{a}\leq j\leq i_{1}<i_{d}\leq j^{\prime}\leq k^{\prime}_{a}\}=\{(x_{[n]\backslash i_{1}},x^{\prime}_{[n]\backslash i_{d}})\}$.
For $\Lambda^{R}_{0}(\bm{x},\bm{x}^{\prime})$, let $T^{R}$ be defined according to (5), then we have the following claim.
Claim $\bm{0^{\prime}}$: Suppose $d=d_{\text{H}}(\bm{x},\bm{x}^{\prime})\geq 2$.
  • 1)
    If $|T^{R}\cap[i_{1}+1,i_{d}]|=0$, then $x_{[n]\backslash i_{d}}=x^{\prime}_{[n]\backslash i_{1}}$ and $\Lambda^{R}_{0}=\Lambda^{R}_{0}(\bm{x},\bm{x}^{\prime})=\{(x_{[n]\backslash i_{d}},x^{\prime}_{[n]\backslash i_{1}})\}$. Hence, we have $\Omega^{R}_{0}=\Omega^{R}_{0}(\bm{x},\bm{x}^{\prime})=B^{\text{S}}_{1}(x_{[n]\backslash i_{d}})=B^{\text{S}}_{1}(x^{\prime}_{[n]\backslash i_{1}})$.
  • 2)
    If $|T^{R}\cap[i_{1}+1,i_{d}]|\geq 1$, then $\Lambda^{R}_{0}(\bm{x},\bm{x}^{\prime})=\emptyset$.
Note that $T^{R}\cap[i_{1}+1,i_{d}]=T^{R}(\bm{x},\bm{x}^{\prime})\cap[i_{1}+1,i_{d}]=T^{L}(\bm{x}^{\prime},\bm{x})\cap[i_{1}+1,i_{d}]$ and $(\bm{z},\bm{z}^{\prime})\in\Lambda^{R}_{0}(\bm{x},\bm{x}^{\prime})$ if and only if $(\bm{z}^{\prime},\bm{z})\in\Lambda^{L}_{0}(\bm{x}^{\prime},\bm{x})$. Also note that $\bm{x},\bm{x}^{\prime}\in\Sigma_{q}^{n}$ are arbitrarily chosen. So, Claim $\bm{0^{\prime}}$ can be obtained directly from Claim $\bm{0}$.
III-B For $\Lambda^{L}_{1}(\bm{x},\bm{x}^{\prime})$ and $\Lambda^{R}_{1}(\bm{x},\bm{x}^{\prime})$
We first consider $\Lambda^{L}_{1}(\bm{x},\bm{x}^{\prime})$. By definition, $\Lambda^{L}_{1}(\bm{x},\bm{x}^{\prime})$ is the set of all $(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}})$ such that $(j,j^{\prime})\in[n]\times[n]$, $j\leq j^{\prime}$ and $d_{\text{H}}(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}})=1$. Then by Observation 2, we have $|S\cap[1,j-1]|+|T^{L}\cap[j+1,j^{\prime}]|+|S\cap[j^{\prime}+1,n]|=1$, and so there are the following three cases to be considered.
1. $|S\cap[1,j-1]|=1$ and $|T^{L}\cap[j+1,j^{\prime}]|=|S\cap[j^{\prime}+1,n]|=0$.
2. $|T^{L}\cap[j+1,j^{\prime}]|=1$ and $|S\cap[1,j-1]|=|S\cap[j^{\prime}+1,n]|=0$.
3. $|S\cap[j^{\prime}+1,n]|=1$ and $|S\cap[1,j-1]|=|T^{L}\cap[j+1,j^{\prime}]|=0$.
For each $i\in\{1,2,3\}$, let $\Lambda^{L}_{1,i}=\Lambda^{L}_{1,i}(\bm{x},\bm{x}^{\prime})$ be the set of all $(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}})\in\Lambda^{L}_{1}(\bm{x},\bm{x}^{\prime})$, where $(j,j^{\prime})\in[n]\times[n]$ and $j\leq j^{\prime}$, such that the conditions of Case $i$ hold. Clearly,
$\Lambda^{L}_{1}=\bigcup_{i=1}^{3}\Lambda^{L}_{1,i}.$
If $T^{L}\cap[1,i_{1}]\neq\emptyset$, we let
$k_{1}=\max(T^{L}\cap[1,i_{1}]);$ (7)
if $T^{L}\cap[i_{d}+1,n]\neq\emptyset$, we let
$k^{\prime}_{1}=\min(T^{L}\cap[i_{d}+1,n]).$ (8)
Then
$2\leq k_{1}\leq i_{1}<i_{d}<k^{\prime}_{1}\leq n$
and we have the following Claims $1.1-1.3$.
Claim 1.1: If $T^{L}\cap[i_{2}+1,i_{d}]\neq\emptyset$, then $\Lambda^{L}_{1,1}(\bm{x},\bm{x}^{\prime})=\emptyset$; if $T^{L}\cap[i_{2}+1,i_{d}]=\emptyset$, then $\Lambda^{L}_{1,1}=\{(x_{[n]\backslash i_{2}},x^{\prime}_{[n]\backslash i_{d}})\}$.
Claim 1.2: If $|T^{L}\cap[i_{1}+1,i_{d}]|\geq 2$, then $\Lambda^{L}_{1,2}(\bm{x},\bm{x}^{\prime})=\emptyset$; if $|T^{L}\cap[i_{1}+1,i_{d}]|=1$, then $\Lambda^{L}_{1,2}=\{(x_{[n]\backslash i_{1}},x^{\prime}_{[n]\backslash i_{d}})\}$; if $|T^{L}\cap[i_{1}+1,i_{d}]|=0$, then we have $\Lambda^{L}_{1,2}\subseteq\{(x_{[n]\backslash k_{1}-1},x^{\prime}_{[n]\backslash i_{d}}),(x_{[n]\backslash i_{1}},x^{\prime}_{[n]\backslash k^{\prime}_{1}})\}$. [^1] Here $\Lambda^{L}_{1,2}(\bm{x},\bm{x}^{\prime})\subseteq\{(x_{[n]\backslash k_{1}-1},x^{\prime}_{[n]\backslash i_{d}}),(x_{[n]\backslash i_{1}},x^{\prime}_{[n]\backslash k^{\prime}_{1}})\}$ means that if $k_{1}~{}($resp. $k^{\prime}_{1})$ exists, then $(x_{[n]\backslash k_{1}-1},x^{\prime}_{[n]\backslash i_{d}})\in\Lambda^{L}_{1,2}(\bm{x},\bm{x}^{\prime})~{}\big{(}$resp. $(x_{[n]\backslash i_{1}},x^{\prime}_{[n]\backslash k^{\prime}_{1}})\in\Lambda^{L}_{1,2}(\bm{x},\bm{x}^{\prime})\big{)}$. The usage of the notation $\subseteq$ in Claims 1.2, 2.2, 2.4, 2.6, 2.2, 2.4 and 2.6 should be understood similarly.
Claim 1.3: If $T^{L}\cap[i_{1}+1,i_{d-1}]\neq\emptyset$, then $\Lambda^{L}_{1,3}(\bm{x},\bm{x}^{\prime})=\emptyset$; if $T^{L}\cap[i_{1}+1,i_{d-1}]=\emptyset$, then $\Lambda^{L}_{1,3}=\{(x_{[n]\backslash i_{1}},x^{\prime}_{[n]\backslash i_{d-1}})\}$.
Similarly, we can divide $\Lambda^{R}_{1}=\Lambda^{R}_{1}(\bm{x},\bm{x}^{\prime})$ into three subsets $\Lambda^{R}_{1,i}=\Lambda^{R}_{1,i}(\bm{x},\bm{x}^{\prime})$, $i=1,2,3~{}($see Table 1), and we can obtain
$\Lambda^{R}_{1}=\bigcup_{i=1}^{3}\Lambda^{R}_{1,i}.$
If $T^{R}\cap[1,i_{1}]\neq\emptyset$, let
$m_{1}=\max(T^{R}\cap[1,i_{1}]);$ (9)
if $T^{R}\cap[i_{d}+1,n]\neq\emptyset$, let
$m^{\prime}_{1}=\min(T^{R}\cap[i_{d}+1,n]).$ (10)
Then
$2\leq m_{1}\leq i_{1}<i_{d}<m^{\prime}_{1}\leq n$
and we have the following Claims ${}^{\prime}-$1.3.
Claim 1.1: If $T^{R}\cap[i_{2}+1,i_{d}]\neq\emptyset$, then $\Lambda^{R}_{1,1}(\bm{x},\bm{x}^{\prime})=\emptyset$; if $T^{R}\cap[i_{2}+1,i_{d}]=\emptyset$, then $\Lambda^{R}_{1,1}=\{(x_{[n]\backslash i_{d}},x^{\prime}_{[n]\backslash i_{2}})\}$.
Claim 1.2: If $|T^{R}\cap[i_{1}+1,i_{d}]|\geq 2$, then $\Lambda^{R}_{1,2}(\bm{x},\bm{x}^{\prime})=\emptyset$; if $|T^{R}\cap[i_{1}+1,i_{d}]|=1$, then $\Lambda^{R}_{1,2}=\{(x_{[n]\backslash i_{d}},x^{\prime}_{[n]\backslash i_{1}})\}$; if $|T^{R}\cap[i_{1}+1,i_{d}]|=0$, then we have $\Lambda^{R}_{1,2}(\bm{x},\bm{x}^{\prime})\subseteq\{(x_{[n]\backslash i_{d},x^{\prime}_{[n]\backslash m_{1}-1}}),(x_{[n]\backslash m^{\prime}_{1}},x^{\prime}_{[n]\backslash i_{1}})\}$.
Claim 1.3: If $T^{R}\cap[i_{1}+1,i_{d-1}]\neq\emptyset$, then $\Lambda^{R}_{1,3}(\bm{x},\bm{x}^{\prime})=\emptyset$; if $T^{R}\cap[i_{1}+1,i_{d-1}]=\emptyset$, then $\Lambda^{R}_{1,3}=\{(x_{[n]\backslash i_{d-1}},x^{\prime}_{[n]\backslash i_{1}})\}$.
$|S\cap[1,j-1]|$ $|T^{X}\cap[j+1,j^{\prime}]|$ $|S\cap[j^{\prime}+1,n]|$
$\Lambda^{X}_{1,1}$ $1$ 00 00
$\Lambda^{X}_{1,2}$ 00 $1$ 00
$\Lambda^{X}_{1,3}$ 00 00 $1$
$\Lambda^{X}_{2,1}$ $2$ 00 00
$\Lambda^{X}_{2,2}$ 00 $2$ 00
$\Lambda^{X}_{2,3}$ 00 00 $2$
$\Lambda^{X}_{2,4}$ 00 $1$ $1$
$\Lambda^{X}_{2,5}$ $1$ 00 $1$
$\Lambda^{X}_{2,6}$ $1$ $1$ 00
Table 1. For each $X\in\{L,R\}$, the set $\Lambda^{X}_{1}$ is divided into three subsets $\Lambda^{X}_{1,i}$, $i=1,2,3$, and the set $\Lambda^{X}_{2}$ is divided into six subsets $\Lambda^{X}_{2,i}$, $i=1,2,\cdots,6$, according to the value of $(|S\cap[1,j-1]|,|T^{X}\cap[j+1,j^{\prime}]|,|S\cap[j^{\prime}+1,n]|)$.
Remark 4
For each $X\in\{L,R\}$ and $i\in\{1,2,3\}$, let
$\Omega^{X}_{1,i}=\Omega^{X}_{1,i}(\bm{x},\bm{x}^{\prime})\triangleq\bigcup_{(\bm{z},\bm{z}^{\prime})\in\Lambda^{X}_{1,i}}B^{\text{S}}_{1}(\bm{z},\bm{z}^{\prime}).$
Then
$\Omega^{X}_{1}=\bigcup_{i=1}^{3}\Omega^{X}_{1,i}.$
Moreover, we can easily obtain $\Omega^{X}_{1,i}$ from $\Lambda^{X}_{1,i}$ by Remark 3. As an example, consider $\Omega^{R}_{1,2}$ with $|T^{R}\cap[i_{1}+1,i_{d}]|=1$. By Claim 1.2, we have $\Lambda^{R}_{1,2}=\{(x_{[n]\backslash i_{d}},x^{\prime}_{[n]\backslash i_{1}})\}$. Let $\{j_{1}^{\prime}\}=T^{R}\cap[i_{1}+1,i_{d}]$. Then by Remark 3, we can obtain $\Omega^{R}_{1,2}=B^{\text{S}}_{1}(x_{[n]\backslash i_{d}},x^{\prime}_{[n]\backslash i_{1}})=\{\phi^{i_{d}}_{j_{1}^{\prime}-1;a}(\bm{x}):a\in\Sigma_{q}\}=\{\phi^{i_{1}}_{j_{1}^{\prime};a}(\bm{x}^{\prime}):a\in\Sigma_{q}\}$. In particular, $|B^{\text{S}}_{1}(\bm{z},\bm{z}^{\prime})|\leq q$ for each $(\bm{z},\bm{z}^{\prime})\in\Lambda^{X}_{1}$ and each $X\in\{L,R\}$.
In the following, we prove Claims 1.1--1.3. Note that $T^{R}=T^{R}(\bm{x},\bm{x}^{\prime})=T^{L}(\bm{x}^{\prime},\bm{x})$ and $(\bm{z},\bm{z}^{\prime})\in\Lambda^{R}_{1}(\bm{x},\bm{x}^{\prime})$ if and only if $(\bm{z}^{\prime},\bm{z})\in\Lambda^{L}_{1}(\bm{x}^{\prime},\bm{x})$. So, Claims ${}^{\prime}-$1.3 can be obtained directly from Claims 1.1--1.3.
Let $k_{a}=\max(T^{L}\cap[1,i_{1}]\backslash k_{1})$ if $|T^{L}\cap[1,i_{1}]|\geq 2$, and $k_{a}=1$ otherwise; let $k^{\prime}_{a}=\min(T^{L}\cap[i_{d}+1,n]\backslash\{k^{\prime}_{1}\})-1$ if $|T^{L}\cap[i_{d}+1,n]|\geq 2$, and $k^{\prime}_{a}=n$ otherwise. Then we have
$k_{a}<k_{1}\leq i_{1}<i_{d}<k^{\prime}_{1}\leq k^{\prime}_{a}.$ (11)
proof 3 (Proof of Claim 1.1)
By definition, $\Lambda^{L}_{1,1}(\bm{x},\bm{x}^{\prime})\neq\emptyset$ if and only if there exist $(j,j^{\prime})$ satisfying conditions of Case 1 ((((i.e., $|S\cap[1,j-1]|=1$ and $|T^{L}\cap[j+1,j^{\prime}]|=|S\cap[j^{\prime}+1,n]|=0)$.
Suppose $(j,j^{\prime})$ satisfying the conditions of Case 1. Then from $|S\cap[1,j-1]|=1$ and $|S\cap[j^{\prime}+1,n]|=0$, we have
$i_{1}<j\leq i_{2}\leq i_{d}\leq j^{\prime}.$
Combining this with $|T^{L}\cap[j+1,j^{\prime}]|=0$ (condition of Case 1), we obtain $T^{L}\cap[i_{2}+1,i_{d}]\subseteq T^{L}\cap[j+1,j^{\prime}]=\emptyset$. Hence, if $T^{L}\cap[i_{2}+1,i_{d}]\neq\emptyset$, then we have $\Lambda^{L}_{1,1}(\bm{x},\bm{x}^{\prime})=\emptyset$.
Conversely, suppose $T^{L}\cap[i_{2}+1,i_{d}]=\emptyset$. We need to prove that $\Lambda^{L}_{1,1}(\bm{x},\bm{x}^{\prime})=\{(x_{[n]\backslash i_{2}},x^{\prime}_{[n]\backslash i_{d}})\}$. Let $j_{1}=\max(T^{L}\cap[i_{1}+1,i_{2}])$ if $T^{L}\cap[i_{1}+1,i_{2}]\neq\emptyset$, and $j_{1}=i_{1}+1$ otherwise. Then by (11), we have
$i_{1}<j_{1}\leq i_{2}\leq i_{d}<k^{\prime}_{1}.$
It is not hard to verify that $(j,j^{\prime})$ satisfies the conditions of Case 1 ((((i.e., $|S\cap[1,j-1]|=1$ and $|T^{L}\cap[j+1,j^{\prime}]|=|S\cap[j^{\prime}+1,n]|=0)$ if and only if
$j_{1}\leq j\leq i_{2}\leq i_{d}\leq j^{\prime}<k^{\prime}_{1}.$ (12)
Therefore, $\Lambda^{L}_{1,1}(\bm{x},\bm{x}^{\prime})=\{(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}}):j_{1}\leq j\leq i_{2}\leq i_{d}\leq j^{\prime}<k^{\prime}_{1}\}\neq\emptyset$. Note that by (12) and by the definition of $j_{1}$ and $k^{\prime}_{1}$, we can obtain $[j_{1}+1,i_{2}]\cap T^{L}=\emptyset$ and $[i_{d}+1,k^{\prime}_{1}-1]\cap T^{L}=\emptyset.$ Moreover, by the definition of $S$ and $j_{1}$, we can obtain $[j_{1},i_{2}-1]\cap S=\emptyset$ and $[i_{d}+1,k^{\prime}_{1}-1]\cap S=\emptyset.$ Hence, by Lemma 1, $x_{[j_{1},i_{2}]}~{}($resp. $x^{\prime}_{[i_{d},k^{\prime}_{1}-1]})$ is contained in a run of $\bm{x}~{}($resp. $\bm{x}^{\prime})$, which implies that $\Lambda^{L}_{1,1}(\bm{x},\bm{x}^{\prime})=\{(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}}):j_{1}\leq j\leq i_{2}\leq i_{d}\leq j^{\prime}<k^{\prime}_{1}\}=\{(x_{[n]\backslash i_{2}},x^{\prime}_{[n]\backslash i_{d}})\}$.
proof 4 (Proof of Claim 1.2)
By definition, $\Lambda^{L}_{1,2}(\bm{x},\bm{x}^{\prime})\neq\emptyset$ if and only if there exist $(j,j^{\prime})$ satisfying conditions of Case 2 ((((i.e., $|S\cap[1,j-1]|=1$ and $|T^{L}\cap[j+1,j^{\prime}]|=|S\cap[j^{\prime}+1,n]|=0)$.
Suppose $(j,j^{\prime})$ satisfying the conditions of Case 2. By the condition $|S\cap[1,j-1]|=|S\cap[j^{\prime}+1,n]|=0$, we have
$j\leq i_{1}<i_{d}\leq j^{\prime}.$
Combining this with $|T^{L}\cap[j+1,j^{\prime}]|=1$ (condition of Case 2), we have $|T^{L}\cap[i_{1}+1,i_{d}]|\leq|T^{L}\cap[j+1,j^{\prime}]|=1$. Hence, $\Lambda^{L}_{1,2}(\bm{x},\bm{x}^{\prime})=\emptyset$ if $|T^{L}\cap[i_{1}+1,i_{d}]|\geq 2$.
Conversely, suppose $|T^{L}\cap[i_{1}+1,i_{d}]|\leq 1$. We have the following two cases (Cases a--b).
Case a: $|T^{L}\cap[i_{1}+1,i_{d}]|=1$. Then by the conditions of Case 2, and by (11), we have $k_{1}\leq j\leq i_{1}<i_{d}\leq j^{\prime}<k^{\prime}_{1}$. Similar to Claim 1.1, we can prove $x_{[k_{1},i_{1}]}~{}($resp. $x^{\prime}_{[i_{d},k^{\prime}_{1}-1]})$ is contained in a run of $\bm{x}~{}($resp. $\bm{x}^{\prime})$, so we have $\Lambda^{L}_{1,2}(\bm{x},\bm{x}^{\prime})=\{(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}}):k_{1}\leq j\leq i_{1}<i_{d}\leq j^{\prime}<k^{\prime}_{1}\}=\{(x_{[n]\backslash i_{1}},x^{\prime}_{[n]\backslash i_{d}})\}$.
Case b: $|T^{L}\cap[i_{1}+1,i_{d}]|=0$. In this case, by conditions of Case 2, and by (11), we have two possibilities: i) $k_{a}\leq j<k_{1}\leq i_{1}<i_{d}\leq j^{\prime}<k^{\prime}_{1}$, which implies $(x_{[n]\backslash k_{1}-1},x^{\prime}_{[n]\backslash i_{d}})\in\Lambda^{L}_{1,2}(\bm{x},\bm{x}^{\prime})$; and ii) $k_{1}\leq j\leq i_{1}<i_{d}<k^{\prime}_{1}\leq j^{\prime}\leq k^{\prime}_{a}$, which implies $(x_{[n]\backslash i_{1}},x^{\prime}_{[n]\backslash k^{\prime}_{1}})\in\Lambda^{L}_{1,2}(\bm{x},\bm{x}^{\prime})$. Thus, we have $\Lambda^{L}_{1,2}(\bm{x},\bm{x}^{\prime})\subseteq\{(x_{[n]\backslash k_{1}-1},x^{\prime}_{[n]\backslash i_{d}}),(x_{[n]\backslash i_{1}},x^{\prime}_{[n]\backslash k^{\prime}_{1}})\}$.
proof 5 (Proof of Claim 1.3)
The proof is similar to Claim 1.1.
First suppose $(j,j^{\prime})$ satisfies conditions of Case 3 ((((i.e., $|S\cap[j^{\prime}+1,n]|=1$ and $|S\cap[1,j-1]|=|T^{L}\cap[j+1,j^{\prime}]|=0)$. Then by $|S\cap[1,j-1]|=0$ and $|S\cap[j^{\prime}+1,n]|=1$, we have
$j\leq i_{1}\leq i_{d-1}\leq j^{\prime}<i_{d}.$
Combining this with $|T^{L}\cap[j+1,j^{\prime}]|=0$, we have $T^{L}\cap[i_{1}+1,i_{d-1}]\subseteq T^{L}\cap[j+1,j^{\prime}]=\emptyset$, which implies that $\Lambda^{L}_{1,3}(\bm{x},\bm{x}^{\prime})=\emptyset$ if $T^{L}\cap[i_{1}+1,i_{d-1}]\neq\emptyset$.
Conversely, suppose $T^{L}\cap[i_{1}+1,i_{d-1}]=\emptyset$. Let $j_{1}^{\prime}=\min(T^{L}\cap[i_{d-1}+1,i_{d}])$ if $T^{L}\cap[i_{d-1}+1,i_{d}]\neq\emptyset$, and $j_{1}^{\prime}=i_{d}$ otherwise. Then by (11), we have
$k_{1}\leq i_{1}\leq i_{d-1}<j^{\prime}_{1}\leq i_{d}.$
Clearly, $(j,j^{\prime})$ satisfies the conditions of Case 3 ((((i.e., $|S\cap[j^{\prime}+1,n]|=1$ and $|S\cap[1,j-1]|=|T^{L}\cap[j+1,j^{\prime}]|=0)$ if and only if
$k_{1}\leq j\leq i_{1}\leq i_{d-1}\leq j^{\prime}<j^{\prime}_{1}.$ (13)
By Lemma 1, we can prove $x_{[k_{1},i_{1}]}~{}($resp. $x^{\prime}_{[i_{d-1},j^{\prime}_{1}-1]})$ is contained in a run of $\bm{x}~{}($resp. $\bm{x}^{\prime})$, which implies that $\Lambda^{L}_{1,3}(\bm{x},\bm{x}^{\prime})=\{(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}}):k_{1}\leq j\leq i_{1}\leq i_{d-1}\leq j^{\prime}<j^{\prime}_{1}\}=\{(x_{[n]\backslash i_{1}},x^{\prime}_{[n]\backslash i_{d-1}})\}$.
III-C For $\Lambda^{L}_{2}(\bm{x},\bm{x}^{\prime})$ and $\Lambda^{R}_{2}(\bm{x},\bm{x}^{\prime})$
We first consider $\Lambda^{L}_{2}(\bm{x},\bm{x}^{\prime})$. Recall that $\Lambda^{L}_{2}(\bm{x},\bm{x}^{\prime})$ is the set of all $(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}})$ such that $(j,j^{\prime})\in[n]\times[n]$, $j\leq j^{\prime}$ and $d_{\text{H}}(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}})=2$. By Observation 2, we have $|S\cap[1,j-1]|+|T^{L}\cap[j+1,j^{\prime}]|+|S\cap[j^{\prime}+1,n]|=2$, and we need to consider the following six cases.
1. $|S\cap[1,j-1]|=2$ and $|T^{L}\cap[j+1,j^{\prime}]|=|S\cap[j^{\prime}+1,n]|=0$.
2. $|T^{L}\cap[j+1,j^{\prime}]|=2$ and $|S\cap[1,j-1]|=|S\cap[j^{\prime}+1,n]|=0$.
3. $|S\cap[j^{\prime}+1,n]|=2$ and $|S\cap[1,j-1]|=|T^{L}\cap[j+1,j^{\prime}]|=0$.
4. $|S\cap[1,j-1]|=0$ and $|T^{L}\cap[j+1,j^{\prime}]|=|S\cap[j^{\prime}+1,n]|=1$.
5. $|T^{L}\cap[j+1,j^{\prime}]|=0$ and $|S\cap[1,j-1]|=|S\cap[j^{\prime}+1,n]|=1$.
6. $|S\cap[j^{\prime}+1,n]|=0$ and $|S\cap[1,j-1]|=|T^{L}\cap[j+1,j^{\prime}]|=1$.
For each $i\in\{1,2,\cdots,6\}$, let $\Lambda^{L}_{2,i}=\Lambda^{L}_{2,i}(\bm{x},\bm{x}^{\prime})$ be the set of $(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}})\in\Lambda^{L}_{2}(\bm{x},\bm{x}^{\prime})$ such that $(j,j^{\prime})\in[n]\times[n]$, $j\leq j^{\prime}$ and conditions of Case $i$ hold. Then
$\Lambda^{L}_{2}=\bigcup_{i=1}^{6}\Lambda^{L}_{2,i}.$
If $|T^{L}\cap[1,i_{1}]|\geq 2$, let
$k_{2}=\max(T^{R}\cap[1,i_{1}]\backslash k_{1})$ (14)
where $k_{1}$ is defined as in (7); if $|T^{R}\cap[i_{d}+1,n]|\geq 2$, let
$k^{\prime}_{2}=\min(T^{R}\cap[i_{d}+1,n]\backslash k^{\prime}_{1})$ (15)
where $k^{\prime}_{1}$ is defined as in (8). Then
$2\leq k_{2}<k_{1}\leq i_{1}<i_{d}<k^{\prime}_{1}<k^{\prime}_{2}\leq n$
and we have the following Claims 2.1--2.6.
Claim 2.1: If $d=2$, then $\Lambda^{L}_{2,1}=\{(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j}):j\in[i_{2}+1,n]\}$ and $|\Lambda^{L}_{2,1}|$ equals to the number of runs of $x_{[i_{2}+1,n]}$; if $d\geq 3$ and $T^{L}\cap[i_{3}+1,i_{d}]\neq\emptyset$, then $\Lambda^{L}_{2,1}=\emptyset$; if $d\geq 3$ and $T^{L}\cap[i_{3}+1,i_{d}]=\emptyset$, then $\Lambda^{L}_{2,1}=\{(x_{[n]\backslash i_{3}},x^{\prime}_{[n]\backslash i_{d}})\}$.
Claim 2.2: If $|T^{L}\cap[i_{1}+1,i_{d}]|\geq 3$, then we have $\Lambda^{L}_{2,2}(\bm{x},\bm{x}^{\prime})=\emptyset$; if $|T^{L}\cap[i_{1}+1,i_{d}]|=2$, then $\Lambda^{L}_{2,2}=\{(x_{[n]\backslash i_{1}},x^{\prime}_{[n]\backslash i_{d}})\}$; if $|T^{L}\cap[i_{1}+1,i_{d}]|=1$, then $\Lambda^{L}_{2,2}\subseteq\{(x_{[n]\backslash k_{1}-1},x^{\prime}_{[n]\backslash i_{d}}),(x_{[n]\backslash i_{1}},x^{\prime}_{[n]\backslash k^{\prime}_{1}})\}$; if $|T^{L}\cap[i_{1}+1,i_{d}]|=0$, then $\Lambda^{L}_{2,2}\subseteq\{(x_{[n]\backslash k_{2}-1},x^{\prime}_{[n]\backslash i_{d}}),(x_{[n]\backslash k_{1}-1},x^{\prime}_{[n]\backslash k^{\prime}_{1}}),(x_{[n]\backslash i_{1}},x^{\prime}_{[n]\backslash k^{\prime}_{2}})\}$.
Claim 2.3: If $d=2$, then $\Lambda^{L}_{2,3}=\{(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j}):j\in[1,i_{1}-1]\}$ and $|\Lambda^{L}_{2,3}|$ equals to the number of runs of $x_{[1,i_{1}-1]}$; if $d\geq 3$ and $T^{L}\cap[i_{1}+1,i_{d-2}]\neq\emptyset$, then $\Lambda^{L}_{2,3}=\emptyset$; if $d\geq 3$ and $T^{L}\cap[i_{1}+1,i_{d-2}]=\emptyset$, then $\Lambda^{L}_{2,3}=\{(x_{[n]\backslash i_{1}},x^{\prime}_{[n]\backslash i_{d-2}})\}$.
Claim 2.4: If $|T^{L}\cap[i_{1}+1,i_{d-1}]|\geq 2$, then $\Lambda^{L}_{2,4}=\emptyset$; if $|T^{L}\cap[i_{1}+1,i_{d-1}]|=1$, then $\Lambda^{L}_{2,4}=\{(x_{[n]\backslash i_{1}},x^{\prime}_{[n]\backslash i_{d-1}})\}$; if $|T^{L}\cap[i_{1}+1,i_{d-1}]|=0$, then $\Lambda^{L}_{2,4}\subseteq\{(x_{[n]\backslash k_{1}-1},x^{\prime}_{[n]\backslash i_{d-1}}),(x_{[n]\backslash i_{1}},x^{\prime}_{[n]\backslash j_{1}^{\prime}})\}$, where $j_{1}^{\prime}=\min(T^{L}\cap[i_{d-1}+1,i_{d}-1])$ when $T^{L}\cap[i_{d-1}+1,i_{d}-1]\neq\emptyset$.
Claim 2.5: If $d=2$, then $\Lambda^{L}_{2,5}=\{(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j}):j\in[i_{1}+1,i_{2}-1]\}$ and $|\Lambda^{L}_{2,5}|$ equals to the number of runs of $x_{[i_{1}+1,i_{2}-1]}$; if $d\geq 3$ and $T^{L}\cap[i_{2}+1,i_{d-1}]\neq\emptyset$, then $\Lambda^{L}_{2,5}=\emptyset$; if $d\geq 3$ and $T^{L}\cap[i_{2}+1,i_{d-1}]=\emptyset$, then $\Lambda^{L}_{2,5}=\{(x_{[n]\backslash i_{2}},x^{\prime}_{[n]\backslash i_{d-1}})\}$.
Claim 2.6: If $|T^{L}\cap[i_{2}+1,i_{d}]|\geq 2$, then $\Lambda^{L}_{2,6}=\emptyset$; if $|T^{L}\cap[i_{2}+1,i_{d}]|=1$, then $\Lambda^{L}_{2,6}=\{(x_{[n]\backslash i_{2}},x^{\prime}_{[n]\backslash i_{d}})\}$; if $|T^{L}\cap[i_{2}+1,i_{d}]|=0$, then $\Lambda^{L}_{2,6}\subseteq\{(x_{[n]\backslash i_{2}},x^{\prime}_{[n]\backslash k^{\prime}_{1}}),(x_{[n]\backslash j_{1}-1},x^{\prime}_{[n]\backslash i_{d}})\}$, where $j_{1}=\max(T^{L}\cap[i_{1}+2,i_{2}])$ when $T^{L}\cap[i_{1}+2,i_{2}]\neq\emptyset$.
Similarly, the set $\Lambda^{R}_{2}=\Lambda^{R}_{2}(\bm{x},\bm{x}^{\prime})$ can be divided into six subsets $\Lambda^{R}_{2,i}=\Lambda^{R}_{2,i}(\bm{x},\bm{x}^{\prime})$, $i=1,2,\cdots,6~{}($see table 1)))), and
$\Lambda^{R}_{2}=\bigcup_{i=1}^{6}\Lambda^{R}_{2,i}.$
If $|T^{L}\cap[1,i_{1}]|\geq 2$, let
$m_{2}=\max(T^{R}\cap[1,i_{1}]\backslash m_{1})$ (16)
where $m_{1}$ is defined as in (9); if $|T^{R}\cap[i_{d}+1,n]|\geq 2$, let
$m^{\prime}_{2}=\min(T^{R}\cap[i_{d}+1,n]\backslash m^{\prime}_{1})$ (17)
where $m^{\prime}_{1}$ is defined as in (10). Then
$2\leq m_{2}<m_{1}\leq i_{1}<i_{d}<m^{\prime}_{1}<m^{\prime}_{2}\leq n$
and we have the following Claims ${}^{\prime}-$2.6.
Claim 2.1: If $d=2$, then $\Lambda^{R}_{2,1}=\{(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j}):j\in[i_{2}+1,n]\}$ and $|\Lambda^{R}_{2,1}|$ equals to the number of runs of $x_{[i_{2}+1,n]}$; if $d\geq 3$ and $T^{R}\cap[i_{3}+1,i_{d}]\neq\emptyset$, then $\Lambda^{R}_{2,1}=\emptyset$; if $d\geq 3$ and $T^{R}\cap[i_{3}+1,i_{d}]=\emptyset$, then $\Lambda^{R}_{2,1}=\{(x_{[n]\backslash i_{d}},x^{\prime}_{[n]\backslash i_{3}})\}$.
Claim 2.2: If $|T^{R}\cap[i_{1}+1,i_{d}]|\geq 3$, then we have $\Lambda^{R}_{2,2}=\emptyset$; if $|T^{R}\cap[i_{1}+1,i_{d}]|=2$, then $\Lambda^{R}_{2,2}=\{(x_{[n]\backslash i_{d}},x^{\prime}_{[n]\backslash i_{1}})\}$; if $|T^{R}\cap[i_{1}+1,i_{d}]|=1$, then $\Lambda^{R}_{2,2}\subseteq\{(x_{[n]\backslash i_{d}},x^{\prime}_{[n]\backslash m_{1}-1}),(x_{[n]\backslash m^{\prime}_{1}},x^{\prime}_{[n]\backslash i_{1}})\}$; if $|T^{R}\cap[i_{1}+1,i_{d}]|=0$, then $\Lambda^{R}_{2,2}\subseteq\{(x_{[n]\backslash i_{d}},x^{\prime}_{[n]\backslash m_{2}-1}),(x_{[n]\backslash m^{\prime}_{1}},x^{\prime}_{[n]\backslash m_{1}-1}),(x_{[n]\backslash m^{\prime}_{2}},x^{\prime}_{[n]\backslash i_{1}})\}$.
Claim 2.3: If $d=2$, then $\Lambda^{R}_{2,3}=\{(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j}):j\in[1,i_{1}-1]\}$ and $|\Lambda^{R}_{2,3}|$ equals to the number of runs of $x_{[1,i_{1}-1]}$; if $d\geq 3$ and $T^{R}\cap[i_{1}+1,i_{d-2}]\neq\emptyset$, then $\Lambda^{R}_{2,3}=\emptyset$; if $d\geq 3$ and $T^{R}\cap[i_{1}+1,i_{d-2}]=\emptyset$, then $\Lambda^{R}_{2,3}=\{(x_{[n]\backslash i_{d-2}},x^{\prime}_{[n]\backslash i_{1}})\}$.
Claim 2.4: If $|T^{R}\cap[i_{1}+1,i_{d-1}]|\geq 2$, then $\Lambda^{R}_{2,4}=\emptyset$; if $|T^{R}\cap[i_{1}+1,i_{d-1}]|=1$, then $\Lambda^{R}_{2,4}=\{(x_{[n]\backslash i_{d-1}},x^{\prime}_{[n]\backslash i_{1}})\}$; if $|T^{R}\cap[i_{1}+1,i_{d-1}]|=0$, then $\Lambda^{R}_{2,4}\subseteq\{(x_{[n]\backslash i_{d-1}},x^{\prime}_{[n]\backslash m_{1}-1}),(x_{[n]\backslash\bar{j}_{1}^{\prime}},x^{\prime}_{[n]\backslash i_{1}})\}$, where $\bar{j}_{1}^{\prime}=\min(T^{R}\cap[i_{d-1}+1,i_{d}-1])$ when $T^{R}\cap[i_{d-1}+1,i_{d}-1]\neq\emptyset$.
Claim 2.5: If $d=2$, then $\Lambda^{R}_{2,5}=\{(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j}):j\in[i_{1}+1,i_{2}-1]\}$ and $|\Lambda^{R}_{2,5}|$ equals to the number of runs of $x_{[i_{1}+1,i_{2}-1]}$; if $d\geq 3$ and $T^{R}\cap[i_{2}+1,i_{d-1}]\neq\emptyset$, then $\Lambda^{R}_{2,5}=\emptyset$; if $d\geq 3$ and $T^{R}\cap[i_{2}+1,i_{d-1}]=\emptyset$, then $\Lambda^{R}_{2,5}=\{(x_{[n]\backslash i_{d-1}},x^{\prime}_{[n]\backslash i_{2}})\}$.
Claim 2.6: If $|T^{R}\cap[i_{2}+1,i_{d}]|\geq 2$, then $\Lambda^{R}_{2,6}=\emptyset$; if $|T^{R}\cap[i_{2}+1,i_{d}]|=1$, then $\Lambda^{R}_{2,6}=\{(x_{[n]\backslash i_{d}},x^{\prime}_{[n]\backslash i_{2}})\}$; if $|T^{R}\cap[i_{2}+1,i_{d}]|=0$, then $\Lambda^{R}_{2,6}\subseteq\{(x_{[n]\backslash m^{\prime}_{1}},x^{\prime}_{[n]\backslash i_{2}}),(x_{[n]\backslash i_{d},x^{\prime}_{[n]\backslash\bar{j}_{1}-1}})\}$, where $\bar{j}_{1}=\max(T^{R}\cap[i_{1}+2,i_{2}])$ when $T^{R}\cap[i_{1}+2,i_{2}]\neq\emptyset$.
Remark 5
For each $X\in\{L,R\}$ and $i\in\{1,2,\cdots,6\}$, let
$\Omega^{X}_{2,i}=\Omega^{X}_{2,i}(\bm{x},\bm{x}^{\prime})\triangleq\bigcup_{(\bm{z},\bm{z}^{\prime})\in\Lambda^{X}_{2,i}}B^{\text{S}}_{1}(\bm{z},\bm{z}^{\prime}).$
Then
$\Omega^{X}_{2}=\bigcup_{i=1}^{6}\Omega^{X}_{2,i}.$
Moreover, we can easily obtain $\Omega^{X}_{2,i}$ from $\Lambda^{X}_{2,i}$ by Remark 2. As an example, consider $\Omega^{L}_{2,4}$ with the assumption of $|T^{L}\cap[i_{1}+1,i_{d-1}]|=0$. By Claim 2.4, we have $\Lambda^{L}_{2,4}\subseteq\{(x_{[n]\backslash k_{1}-1},x^{\prime}_{[n]\backslash i_{d-1}}),(x_{[n]\backslash i_{1}},x^{\prime}_{[n]\backslash j_{1}^{\prime}})\}$, where $j_{1}^{\prime}=\min(T^{L}\cap[i_{d-1}+1,i_{d}-1])$ when $T^{L}\cap[i_{d-1}+1,i_{d}-1]\neq\emptyset$. Then $\Omega^{L}_{2,4}=B^{\text{S}}_{1}(x_{[n]\backslash k_{1}-1},x^{\prime}_{[n]\backslash i_{d-1}})\cup B^{\text{S}}_{1}(x_{[n]\backslash i_{1}},x^{\prime}_{[n]\backslash j_{1}^{\prime}})$ and by Remark 2, we have $B^{\text{S}}_{1}(x_{[n]\backslash k_{1}-1},x^{\prime}_{[n]\backslash i_{d-1}})=\{\phi^{k_{1}-1}_{k_{1};x^{\prime}_{k_{1}-1}}(\bm{x}),\phi^{k_{1}-1}_{i_{d};x^{\prime}_{i_{d}}}(\bm{x})\}=\{\phi^{i_{d-1}}_{k_{1}-1;x_{k_{1}}}(\bm{x}^{\prime}),\phi^{i_{d-1}}_{i_{d};x_{i_{d}}}(\bm{x}^{\prime})\}$ and $B^{\text{S}}_{1}(x_{[n]\backslash i_{1}},x^{\prime}_{[n]\backslash j_{1}^{\prime}})=\{\phi^{i_{1}}_{i_{1};x^{\prime}_{j_{1}^{\prime}-1}}(\bm{x}),\phi^{i_{1}}_{i_{d};x^{\prime}_{i_{d}}}(\bm{x})\}=\{\phi^{j_{1}^{\prime}}_{j_{1}^{\prime}-1;x_{j_{1}^{\prime}}}(\bm{x}^{\prime}),\phi^{j_{1}^{\prime}}_{i_{d};x_{i_{d}}}(\bm{x}^{\prime})\}$. In particular, $|B^{\text{S}}_{1}(\bm{z},\bm{z}^{\prime})|\leq 2$ for each $(\bm{z},\bm{z}^{\prime})\in\Lambda^{X}_{2}$ and each $X\in\{L,R\}$.
In the following, we prove Claims 2.1--2.6. Note that $T^{R}=T^{R}(\bm{x},\bm{x}^{\prime})=T^{L}(\bm{x}^{\prime},\bm{x})$ and $(\bm{z},\bm{z}^{\prime})\in\Lambda^{R}_{\ell}(\bm{x},\bm{x}^{\prime})$ if and only if $(\bm{z}^{\prime},\bm{z})\in\Lambda^{L}_{\ell}(\bm{x}^{\prime},\bm{x})$. So, Claims ${}^{\prime}-$2.6 can be obtained from Claims 2.1--2.6 directly.
Let $k_{a}=\max(T^{L}\cap[1,i_{1}]\backslash\{k_{1},k_{2}\})$ if $|T^{L}\cap[1,i_{1}]|\geq 3$, and $k_{a}=1$ otherwise, where $k_{1}$ and $k_{2}$ are defined by (7) and (14), respectively. Let $k^{\prime}_{a}=\min(T^{L}\cap[i_{d}+1,n]\backslash\{k^{\prime}_{1},k^{\prime}_{2}\})-1$ if $|T^{L}\cap[i_{d}+1,n]|\geq 3$, and $k_{a}^{\prime}=n$ otherwise, where $k_{1}^{\prime}$ and $k^{\prime}_{2}$ are defined by (8) and (15), respectively. Then we have
$k_{a}<k_{2}<k_{1}\leq i_{1}<i_{d}<k^{\prime}_{1}<k^{\prime}_{2}\leq k^{\prime}_{a}.$ (18)
proof 6 (Proof of Claim 2.1)
Note that $\Lambda^{L}_{2,1}(\bm{x},\bm{x}^{\prime})\neq\emptyset$ if and only if there exists $(j,j^{\prime})$ satisfying the conditions of Case 1. The proof is similar to Claim 1.1.
For $d=2$, it is easy to see that $(j,j^{\prime})$ satisfies the conditions of Case 1 ((((that is, $|S\cap[1,j-1]|=2$ and $|T^{L}\cap[j+1,j^{\prime}]|=|S\cap[j^{\prime}+1,n]|=0)$ if and only if $i_{2}<j\leq j^{\prime}$ and $|T^{L}\cap[j+1,j^{\prime}]|=0$. By definition of $S$ and by Lemma 1, we can prove that $x_{[j,j^{\prime}]}=x^{\prime}_{[j,j^{\prime}]}$ is contained in a run of $x_{[i_{2}+1,n]}=x^{\prime}_{[i_{2}+1,n]}$, so $\Lambda^{L}_{2,1}=\{(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j}):j\in[i_{2}+1,n]\}$ and $|\Lambda^{L}_{2,1}|$ equals to the number of runs of $x_{[i_{2}+1,n]}$.
Suppose $d\geq 3$ and $(j,j^{\prime})$ satisfy the conditions of Case 1. By $|S\cap[1,j-1]|=2$ and $|S\cap[j^{\prime}+1,n]|=0$, we have
$i_{2}<j\leq i_{3}\leq i_{d}\leq j^{\prime}.$
Combining this with $|T^{L}\cap[j+1,j^{\prime}]|=0$, we can obtain $T^{L}\cap[i_{3}+1,i_{d}]\subseteq T^{L}\cap[j+1,j^{\prime}]=\emptyset$, which implies that if $|T^{L}\cap[i_{3}+1,i_{d}]|\neq\emptyset$, then $\Lambda^{L}_{2,1}(\bm{x},\bm{x}^{\prime})=\emptyset$.
Conversely, suppose $d\geq 3$ and $T^{L}\cap[i_{3}+1,i_{d}]=\emptyset$. We need to prove $\Lambda^{L}_{2,1}=\{(x_{[n]\backslash i_{3}},x^{\prime}_{[n]\backslash i_{d}})\}$. Let $j_{1}=\max(T^{L}\cap[i_{2}+1,i_{3}])$ if $T^{L}\cap[i_{2}+1,i_{3}]\neq\emptyset$, and $j_{1}=i_{2}+1$ otherwise. Then by (18), we have
$i_{2}<j_{1}\leq i_{3}\leq i_{d}<k^{\prime}_{1}.$
Clearly, $(j,j^{\prime})$ satisfies the conditions of Case 1 ((((i.e., $|S\cap[1,j-1]|=2$ and $|T^{L}\cap[j+1,j^{\prime}]|=|S\cap[j^{\prime}+1,n]|=0)$ if and only if it satisfies
$i_{2}<j_{1}\leq j\leq i_{3}\leq i_{d}\leq j^{\prime}<k^{\prime}_{1}.$
Moreover, by the definition of $j_{1}$ and $k^{\prime}_{1}$, we have
$[j_{1}+1,i_{3}]\cap T=[i_{d}+1,k^{\prime}_{1}-1]\cap T=\emptyset,$
and
$[j_{1},i_{3}-1]\cap S=[i_{d}+1,k^{\prime}_{1}-1]\cap S=\emptyset.$
Hence, by Lemma 1, $x_{[j_{1},i_{3}]}~{}($resp. $x^{\prime}_{[i_{d},k^{\prime}_{1}-1]})$ is contained in a run of $\bm{x}~{}($resp. $\bm{x}^{\prime})$, which implies that $\Lambda^{L}_{2,1}=\{(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}}):i_{2}<j_{1}\leq j\leq i_{3}\leq i_{d}\leq j^{\prime}<k^{\prime}_{1}\}=\{(x_{[n]\backslash i_{3}},x^{\prime}_{[n]\backslash i_{d}})\}$.
proof 7 (Proof of Claim 2.2)
The proof is similar to Claim 1.2.
Suppose there exists $(j,j^{\prime})$ satisfying the conditions of Case 2 ((((i.e., $|T^{L}\cap[j+1,j^{\prime}]|=2$ and $|S\cap[1,j-1]|=|S\cap[j^{\prime}+1,n]|=0)$. By $|S\cap[1,j-1]|=|S\cap[j^{\prime}+1,n]|=0$, we have
$j\leq i_{1}<i_{d}\leq j^{\prime}.$
Combining this with $|T^{L}\cap[j+1,j^{\prime}]|=2$, we have $|T^{L}\cap[i_{1}+1,i_{d}]|\leq|T^{L}\cap[j+1,j^{\prime}]|=2$, which implies that $\Lambda^{L}_{2,2}(\bm{x},\bm{x}^{\prime})=\emptyset$ if $|T^{L}\cap[i_{1}+1,i_{d}]|\geq 3$.
Conversely, suppose $|T^{L}\cap[i_{1}+1,i_{d}]|\leq 2$. We have the following three cases (Cases a--c).
Case a: $|T^{L}\cap[i_{1}+1,i_{d}]|=2$. By (18), it is easy to see that $(j,j^{\prime})$ satisfying the conditions of Case 2 if and only if $k_{1}\leq j\leq i_{1}<i_{d}\leq j^{\prime}<k^{\prime}_{1}$. Similar to Claim 1.1, we can prove ((((by Lemma 1)))) that $x_{[k_{1},i_{1}]}~{}($resp. $x^{\prime}_{[i_{d},k^{\prime}_{1}-1]})$ is contained in a run of $\bm{x}~{}($resp. $\bm{x}^{\prime})$, which implies that $\Lambda^{L}_{2,2}=\{(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}}):k_{1}\leq j\leq i_{1}<i_{d}\leq j^{\prime}<k^{\prime}_{1}\}=\{(x_{[n]\backslash i_{1}},x^{\prime}_{[n]\backslash i_{d}})\}$.
Case b: $|T^{L}\cap[i_{1}+1,i_{d}]|=1$. By (18), it is easy to see that $(j,j^{\prime})$ satisfying the conditions of Case 2 if and only if it satisfies one of the following two conditions: i) $k_{2}\leq j<k_{1}\leq i_{1}<i_{d}\leq j^{\prime}<k^{\prime}_{1}$; and ii) $k_{1}\leq j\leq i_{1}<i_{3}<k^{\prime}_{1}\leq j^{\prime}<k^{\prime}_{2}$. Hence, we have $\Lambda^{L}_{2,2}\subseteq\{(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}}):\text{condition i) holds}\}\cup\{(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}}):\text{condition ii) holds}\}=\{(x_{[n]\backslash k_{1}-1},x^{\prime}_{[n]\backslash i_{d}}),(x_{[n]\backslash i_{1}},x^{\prime}_{[n]\backslash k^{\prime}_{1}})\}$, where the equality is obtained by applying Lemma 1.
Case c: $|T^{L}\cap[i_{1}+1,i_{d}]|=0$. By (18), it is easy to see that $(j,j^{\prime})$ satisfying the conditions of Case 2 if and only if it satisfies one of the following three conditions: i) $k_{a}\leq j<k_{2}<k_{1}\leq i_{1}<i_{d}\leq j^{\prime}<k^{\prime}_{1}$; ii) $k_{2}\leq j<k_{1}\leq i_{1}<i_{d}<k^{\prime}_{1}\leq j^{\prime}<k^{\prime}_{2}$; and iii) $k_{1}\leq j\leq i_{1}<i_{d}<k_{1}^{\prime}<k^{\prime}_{2}\leq j^{\prime}\leq k^{\prime}_{a}$. Hence, we have $\Lambda^{L}_{2,2}\subseteq\big{\{}(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}}):\text{condition i) holds}\big{\}}\cup\big{\{}(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}}):\text{condition ii) holds}\big{\}}\cup\big{\{}(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}}):\text{condition iii) holds}\big{\}}=\big{\{}(x_{[n]\backslash k_{2}-1},x^{\prime}_{[n]\backslash i_{d}})$, $(x_{[n]\backslash k_{1}-1},x^{\prime}_{[n]\backslash k^{\prime}_{1}})$, $(x_{[n]\backslash i_{1}},x^{\prime}_{[n]\backslash k^{\prime}_{2}})\big{\}}$, where the equality is obtained by applying Lemma 1.
proof 8 (Proof of Claim 2.3)
The proof is similar to Claim 2.1.
For $d=2$, $(j,j^{\prime})$ satisfies the conditions of Case 3 ((((i.e., $|S\cap[j^{\prime}+1,n]|=2$ and $|S\cap[1,j-1]|=|T^{L}\cap[j+1,j^{\prime}]|=0)$ if and only if $j\leq j^{\prime}<i_{1}$ and $|T^{L}\cap[j+1,j^{\prime}]|=0$. By Lemma 1, $x_{[j,j^{\prime}]}=x^{\prime}_{[j,j^{\prime}]}$ is contained in a run of $x_{[1,i_{1}-1]}=x_{[1,i_{1}-1]}$, so $\Lambda^{L}_{2,3}=\{(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j}):j\in[1,i_{1}-1]\}$ and $|\Lambda^{L}_{2,3}|$ equals to the number of runs of $x_{[1,i_{1}-1]}$.
Suppose $d\geq 3$ and $(j,j^{\prime})$ satisfies the conditions of Case 3 ((((i.e., $|S\cap[j^{\prime}+1,n]|=2$ and $|S\cap[1,j-1]|=|T^{L}\cap[j+1,j^{\prime}]|=0)$. By the conditions $|S\cap[j^{\prime}+1,n]|=2$ and $|S\cap[1,j-1]|=|T^{L}\cap[j+1,j^{\prime}]|=0$, we must have $k_{1}\leq j\leq i_{1}\leq i_{d-2}\leq j^{\prime}<j_{1}^{\prime}$ and $T^{L}\cap[i_{1}+1,i_{d-2}]=\emptyset$, where $j_{1}^{\prime}=\min(T^{L}\cap[i_{d-2}+1,i_{d-1}]$ if $T^{L}\cap[i_{d-2}+1,i_{d-1}]\neq\emptyset$, and $j_{1}^{\prime}=i_{d-1}$ otherwise. So, if $T^{L}\cap[i_{1}+1,i_{d-2}]\neq\emptyset$, then $\Lambda^{L}_{2,3}$. Conversely, if $T^{L}\cap[i_{1}+1,i_{d-2}]=\emptyset$, then $(j,j^{\prime})$ satisfies the conditions of Case 3 if and only if $k_{1}\leq j\leq i_{1}\leq i_{d-2}\leq j^{\prime}<j_{1}^{\prime}$. Moreover, by Lemma 1, $x_{[k_{1},i_{1}]}~{}($resp $x^{\prime}_{[i_{d-2},j^{\prime}_{1}-1]})$ is contained in a run of $\bm{x}~{}($resp. $\bm{x}^{\prime})$. Thus, we have $\Lambda^{L}_{2,3}(\bm{x},\bm{x}^{\prime})=\{(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}}):k_{1}\leq j\leq i_{1}\leq i_{d-2}\leq j^{\prime}<j_{1}^{\prime}\}=\{(x_{[n]\backslash i_{1}},x^{\prime}_{[n]\backslash i_{d-2}})\}$.
proof 9 (Proof of Claim 2.4)
Suppose there exists $(j,j^{\prime})$ satisfying the conditions of Case 4 ((((i.e., $|S\cap[1,j-1]|=0$ and $|T^{L}\cap[j+1,j^{\prime}]|=|S\cap[j^{\prime}+1,n]|=1)$. From the conditions $|S\cap[1,j-1]|=0$ and $|S\cap[j^{\prime}+1,n]|=1$, we have
$j\leq i_{1}<i_{d-1}\leq j^{\prime}<i_{d}.$
Combining this with $|T^{L}\cap[j+1,j^{\prime}]|=1$, we have $|T^{L}\cap[i_{1}+1,i_{d-1}]|\leq|T^{L}\cap[j+1,j^{\prime}]|=1$, which implies that if $|T^{L}\cap[i_{1}+1,i_{d-1}]|\geq 2$, then $\Lambda^{L}_{2,4}=\emptyset$.
Suppose $|T^{L}\cap[i_{1}+1,i_{d-1}]|=1$. ((((Note that this condition holds only if $i_{1}<i_{d-1}$, i.e., $d\geq 3.)$ Let $j_{1}^{\prime}=\min(T^{L}\cap[i_{d-1}+1,i_{d}])$ if $T^{L}\cap[i_{d-1}+1,i_{d}]\neq\emptyset$, and $j_{1}^{\prime}=i_{d}$ otherwise. Then $(j,j^{\prime})$ satisfies the conditions of Case 4 if and only if
$k_{1}\leq j\leq i_{1}<i_{d-1}\leq j^{\prime}<j_{1}^{\prime}.$
Moreover, by Lemma 1, we can prove that $x_{[k_{1},i_{1}]}~{}($resp $x^{\prime}_{[i_{d-1},j^{\prime}_{1}-1]})$ is contained in a run of $\bm{x}~{}($resp. $\bm{x}^{\prime})$. Thus, we have $\Lambda^{L}_{2,4}(\bm{x},\bm{x}^{\prime})=\{(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}}):k_{1}\leq j\leq i_{1}<i_{d-1}\leq j^{\prime}<j_{1}^{\prime}\}=\{(x_{[n]\backslash i_{1}},x^{\prime}_{[n]\backslash i_{d-1}})\}$.
Suppose $|T^{L}\cap[i_{1}+1,i_{d-1}]|=0$. We are to prove $\Lambda^{L}_{2,4}\subseteq\{(x_{[n]\backslash k_{1}-1},x^{\prime}_{[n]\backslash i_{d-1}}),(x_{[n]\backslash i_{1}},x^{\prime}_{[n]\backslash j_{1}^{\prime}})\}$, where $j_{1}^{\prime}=\min(T^{L}\cap[i_{d-1}+1,i_{d}-1])$ when $T^{L}\cap[i_{d-1}+1,i_{d}-1]\neq\emptyset$. We need to consider the following Cases a and b.
Case a: $T^{L}\cap[i_{d-1}+1,i_{d}-1]\neq\emptyset$. Let $j_{2}^{\prime}=\min(T^{L}\cap[i_{d-1}+1,i_{d}]\backslash j_{1}^{\prime})$ if $T^{L}\cap[i_{d-1}+1,i_{d}]\neq\emptyset$, and $j_{2}^{\prime}=i_{d}$ otherwise. Then $(j,j^{\prime})$ satisfies the conditions of Case 4 if and only if one of the following two conditions holds: i) $k_{2}\leq j<k_{1}\leq i_{1}\leq i_{d-1}\leq j^{\prime}<j_{1}^{\prime}$; ii) $k_{1}\leq j\leq i_{1}\leq i_{d-1}<j_{1}^{\prime}\leq j^{\prime}<j_{2}^{\prime}$. Hence, we have $\Lambda^{L}_{2,4}\subseteq\{(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}}):\text{condition i) holds}\}\cup\{(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}}):\text{condition ii) holds}\}=\{(x_{[n]\backslash k_{1}-1},x^{\prime}_{[n]\backslash i_{d-1}}),(x_{[n]\backslash i_{1}},x^{\prime}_{[n]\backslash j^{\prime}_{1}})\}$, where the equality is obtained by applying Lemma 1.
Case b: $T^{L}\cap[i_{d-1}+1,i_{d}-1]=\emptyset$. It is easy to see that $(j,j^{\prime})$ satisfies the conditions of Case 4 if and only if $k_{2}\leq j<k_{1}\leq i_{1}\leq i_{d-1}\leq j^{\prime}<j_{1}^{\prime}$. Hence, we have $\Lambda^{L}_{2,4}=\{(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}}):k_{2}\leq j<k_{1}\leq i_{1}\leq i_{d-1}\leq j^{\prime}<j_{1}^{\prime}\}=\{(x_{[n]\backslash k_{1}-1},x^{\prime}_{[n]\backslash i_{d-1}})\}$.
proof 10 (Proof of Claim 2.5)
For $d=2$, $(j,j^{\prime})$ satisfies the conditions of Case 5 ((((i.e., $|S\cap[1,j-1]|=|S\cap[j^{\prime}+1,n]|=1$ and $|T^{L}\cap[j+1,j^{\prime}]|=0)$ if and only if $i_{1}<j\leq j^{\prime}<i_{2}$ and $|T^{L}\cap[j+1,j^{\prime}]|=0$. Similar to Claim 2.1, $x_{[j,j^{\prime}]}=x^{\prime}_{[j,j^{\prime}]}$ is contained in a run of $x_{[i_{1}+1,i_{2}-1]}=x^{\prime}_{[i_{1}+1,i_{2}-1]}$, so $\Lambda^{L}_{2,5}=\{(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j}):j\in[i_{1}+1,i_{2}-1]\}$ and $|\Lambda^{L}_{2,5}|$ equals to the number of runs of $x_{[i_{1}+1,i_{2}-1]}$.
Suppose $d\geq 3$ and $(j,j^{\prime})$ satisfies the conditions of Case 5. From $|S\cap[1,j-1]|=|S\cap[j^{\prime}+1,n]|=1$, we have
$i_{1}<j\leq i_{2}\leq i_{d-1}\leq j^{\prime}<i_{d}.$
So, $T^{L}\cap[i_{2}+1,i_{d-1}]\subseteq T^{L}\cap[j+1,j^{\prime}]=\emptyset$ (the equality is a condition of Case 5), which implies that if $T^{L}\cap[i_{2}+1,i_{d-1}]\neq\emptyset$, then $\Lambda^{L}_{2,5}=\emptyset$.
Now, suppose $d\geq 3$ and $|T^{L}\cap[i_{2}+1,i_{d-1}]|=0$. We are to prove $\Lambda^{L}_{2,5}=\{(x_{[n]\backslash i_{2}},x^{\prime}_{[n]\backslash i_{d-1}})\}$. Let $j_{1}=\max(T^{L}\cap[i_{1}+1,i_{2}])$ if $T^{L}\cap[i_{1}+1,i_{2}\neq\emptyset$, and $j_{1}=i_{1}+1$ otherwise; let $j_{1}^{\prime}=\min(T^{L}\cap[i_{d-1}+1,i_{d}])-1$ if $T^{L}\cap[i_{d-1}+1,i_{d}]\neq\emptyset$, and $j_{1}^{\prime}=i_{d}$ otherwise. Then $(j,j^{\prime})$ satisfies the conditions of Case 5 if and only if it satisfies
$j_{1}\leq j\leq i_{2}\leq i_{d-1}\leq j^{\prime}<j_{1}^{\prime}.$
Moreover, by Lemma 1, we can prove that $x_{[j_{1},i_{2}]}~{}($resp $x^{\prime}_{[i_{d-1},j^{\prime}_{1}-1]})$ is contained in a run of $\bm{x}~{}($resp. $\bm{x}^{\prime})$. Thus, we have $\Lambda^{L}_{2,5}(\bm{x},\bm{x}^{\prime})=\{(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}}):j_{1}\leq j\leq i_{2}\leq i_{d-1}\leq j^{\prime}<j_{1}^{\prime}\}=\{(x_{[n]\backslash i_{2}},x^{\prime}_{[n]\backslash i_{d-1}})\}$.
proof 11 (Proof of Claim 2.6)
The proof is similar to Claim 2.4.
Suppose $(j,j^{\prime})$ satisfies the conditions of Case 6 ((((i.e., $|S\cap[j^{\prime}+1,n]|=0$ and $|S\cap[1,j-1]|=|T^{L}\cap[j+1,j^{\prime}]|=1)$. From the conditions $|S\cap[1,j-1]|=1$ and $|S\cap[j^{\prime}+1,n]|=0$, we have
$i_{1}<j\leq i_{2}<i_{d}\leq j^{\prime}.$
Combining this with $|T^{L}\cap[j+1,j^{\prime}]|=1$, we have $|T^{L}\cap[i_{2}+1,i_{d}]|\leq|T^{L}\cap[j+1,j^{\prime}]|=1$, which implies that if $|T^{L}\cap[i_{2}+1,i_{d}]|\geq 2$, then $\Lambda^{L}_{2,6}=\emptyset$.
Suppose $|T^{L}\cap[i_{2}+1,i_{d}]|=1$. ((((Note that this condition holds only if $i_{2}<i_{d}$, i.e., $d\geq 3.)$ Let $j_{1}=\min(T^{L}\cap[i_{1}+1,i_{2}])$ if $T^{L}\cap[i_{1}+1,i_{2}]\neq\emptyset$, and $j_{1}=i_{d}$ otherwise. Then $(j,j^{\prime})$ satisfies the conditions of Case 6 if and only if
$j_{1}\leq j\leq i_{2}<i_{d}\leq j^{\prime}<k_{1}^{\prime}.$
Moreover, by Lemma 1, we can prove that $x_{[j_{1},i_{2}]}~{}($resp $x^{\prime}_{[i_{d},k^{\prime}_{1}-1]})$ is contained in a run of $\bm{x}~{}($resp. $\bm{x}^{\prime})$. Thus, we have $\Lambda^{L}_{2,6}(\bm{x},\bm{x}^{\prime})=\{(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}}):j_{1}\leq j\leq i_{2}<i_{d}\leq j^{\prime}<k_{1}^{\prime}\}=\{(x_{[n]\backslash i_{2}},x^{\prime}_{[n]\backslash i_{d}})\}$.
Suppose $|T^{L}\cap[i_{2}+1,i_{d}]|=0$. We are to prove $\Lambda^{L}_{2,6}\subseteq\{(x_{[n]\backslash i_{2}},x^{\prime}_{[n]\backslash k^{\prime}_{1}}),(x_{[n]\backslash j_{1}-1},x^{\prime}_{[n]\backslash i_{d}})\}$, where $j_{1}=\max(T^{L}\cap[i_{1}+2,i_{2}])$ when $T^{L}\cap[i_{1}+2,i_{2}]\neq\emptyset$. We need to consider the following Cases a and b.
Case a: $T^{L}\cap[i_{1}+2,i_{2}]\neq\emptyset$. Let $j_{2}=\max(T^{L}\cap[i_{1}+2,i_{2}]\backslash j_{1})$ if $T^{L}\cap[i_{1}+2,i_{2}]\neq\emptyset$, and $j_{2}=i_{1}+1$ otherwise. Then $(j,j^{\prime})$ satisfies the conditions of Case 6 if and only if one of the following two conditions holds: i) $j_{1}\leq j\leq i_{2}\leq i_{d}<k_{1}^{\prime}\leq j^{\prime}<k_{2}^{\prime}$; ii) $j_{2}\leq j<j_{1}\leq i_{2}\leq i_{d}\leq j^{\prime}<k^{\prime}_{1}$. Hence, we have $\Lambda^{L}_{2,4}\subseteq\{(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}}):\text{condition i) holds}\}\cup\{(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}}):\text{condition ii) holds}\}=\{(x_{[n]\backslash i_{2}},x^{\prime}_{[n]\backslash k^{\prime}_{1}}),(x_{[n]\backslash j_{1}-1},x^{\prime}_{[n]\backslash i_{d}})\}$, where the equality is obtained by applying Lemma 1.
Case b: $T^{L}\cap[i_{1}+2,i_{2}]=\emptyset$. It is easy to see that $(j,j^{\prime})$ satisfies the conditions of Case 6 if and only if $j_{1}\leq j\leq i_{2}\leq i_{d}<k_{1}^{\prime}\leq j^{\prime}<k_{2}^{\prime}$. Hence, we have $\Lambda^{L}_{2,4}=\{(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}}):j_{1}\leq j\leq i_{2}\leq i_{d}<k_{1}^{\prime}\leq j^{\prime}<k_{2}^{\prime}\}=\{(x_{[n]\backslash i_{2}},x^{\prime}_{[n]\backslash k^{\prime}_{1}})\}$.
IV Proof of Theorem 1
In this section, we prove Theorem 1. Note that from Claims 1.1--1.3, Claims ${}^{\prime}-$1.3, Claims 2.1--2.6 and Claims ${}^{\prime}-$2.6, we can directly obtain
$\left|B^{\text{D,S}}_{1,1}(\bm{x},\bm{x}^{\prime})\right|$ $=\left|\bigcup_{\ell=0}^{2}(\Omega^{L}_{\ell}\cup\Omega^{R}_{\ell})\right|$
$\leq\sum_{\ell=0}^{2}(|\Omega^{L}_{\ell}|+|\Omega^{R}_{\ell}|)$
$\leq 2(1+(q-1)(n-1))+2(4q)+2(2(n+7))$
$=2(q+1)n+6q+32.$
However, this bound is not tight. To obtain a tight bound of $\big{|}B^{\text{D,S}}_{1,1}(\bm{x},\bm{x}^{\prime})\big{|}$, we need to exclude the intersection of these subsets of $B^{\text{D,S}}_{1,1}(\bm{x},\bm{x}^{\prime})$.
The following remark will be used to exclude repeat count of some sequences in $\bigcup_{\ell=0}^{2}(\Omega^{L}_{\ell}\cup\Omega^{R}_{\ell})$.
Remark 6
If $|T^{L}\cap[i_{1}+1,i_{d}]|=0$, then for any $(\bm{z},\bm{z}^{\prime})\in\Lambda$ such that $\bm{z}=x_{[n]\backslash i_{1}}$ or $\bm{z}^{\prime}=x^{\prime}_{[n]\backslash i_{d}}$, we have $B^{\text{S}}_{1}(\bm{z},\bm{z}^{\prime})\subseteq\Omega^{L}_{0}$; if $|T^{R}\cap[i_{1}+1,i_{d}]|=0$, then for any $(\bm{z},\bm{z}^{\prime})\in\Lambda$ such that $\bm{z}=x_{[n]\backslash i_{d}}$ or $\bm{z}^{\prime}=x^{\prime}_{[n]\backslash i_{1}}$, we have $B^{\text{S}}_{1}(\bm{z},\bm{z}^{\prime})\subseteq\Omega^{R}_{0}$. In fact, by Claim 0, we have $x_{[n]\backslash i_{1}}=x^{\prime}_{[n]\backslash i_{d}}$ and $\Omega^{L}_{0}=B^{\text{S}}_{1}(x_{[n]\backslash i_{1}})=B^{\text{S}}_{1}(x^{\prime}_{[n]\backslash i_{d}})$, so if $\bm{z}=x_{[n]\backslash i_{1}}$ or $\bm{z}^{\prime}=x^{\prime}_{[n]\backslash i_{d}}$, then $B^{\text{S}}_{1}(\bm{z},\bm{z}^{\prime})\subseteq\big{(}B^{\text{S}}_{1}(x_{[n]\backslash i_{1}})\cup B^{\text{S}}_{1}(x^{\prime}_{[n]\backslash i_{d}})\big{)}=\Omega^{L}_{0}$. The other statement can be proved similarly.
To prove the upper bound of $B^{\text{D,S}}_{1,1}(\bm{x},\bm{x}^{\prime})$ in Theorem 1, we need the following Lemmas 2--4.
Lemma 2
For each $X\in\{L,R\}$, the following hold.
  • 1)
    If $|T^{X}\cap[i_{1}+1,i_{d}]|=0$, then we have $\Omega^{X}_{1}\subseteq\Omega^{X}_{0}$.
  • 2)
    If $|T^{X}\cap[i_{1}+1,i_{d}]|\neq 0$ and $d=2$, then $|\Lambda^{L}_{1}|\leq 3$.
  • 3)
    If $|T^{X}\cap[i_{1}+1,i_{d}]|\neq 0$ and $d\geq 3$, then $|\Lambda^{X}_{1}|\leq 2$.
proof 12
We only consider $X=L$. The proof for $X=R$ is similar. By checking Claims 1.1--1.3, we have the following:
1) If $|T^{X}\cap[i_{1}+1,i_{d}]|=0$, then for each $(\bm{z},\bm{z}^{\prime})\in\Lambda_{1}$, either $\bm{z}=x_{[n]\backslash i_{1}}$ or $\bm{z}^{\prime}=x^{\prime}_{[n]\backslash i_{d}}$. Hence, by Remark 6, we have $\Omega^{L}_{1}=\bigcup_{(\bm{z},\bm{z}^{\prime})\in\Lambda^{L}_{1}}B^{\text{S}}_{1}(\bm{z},\bm{z}^{\prime})\subseteq\Omega^{L}_{0}.$
2) If $|T^{X}\cap[i_{1}+1,i_{d}]|\neq 0$ and $d=2$, we have $|\Lambda^{L}_{1,1}|=|\Lambda^{L}_{1,3}|=1$ and $|\Lambda^{L}_{1,2}|\leq 1$. Hence, we have $|\Lambda^{L}_{1}|\leq 3$.
3) If $|T^{X}\cap[i_{1}+1,i_{d}]|\neq 0$ and $d\geq 3$, we have $T^{L}\cap[i_{2}+1,i_{d}]\neq\emptyset$ or $T^{L}\cap[i_{1}+1,i_{d-1}]\neq\emptyset$. Then by Claims 1.1 and 1.3, $|\Lambda^{L}_{1,1}|+|\Lambda^{L}_{1,3}|\leq 1$. Moreover, by Claim 1.2, $|\Lambda^{L}_{1,2}|\leq 1$. Thus, we have $|\Lambda^{L}_{1}|=|\bigcup_{i=1}^{3}\Lambda^{L}_{1,i}|\leq 2$.
For $d=2$, to simplify the expressions, we need the following notations. For $X\in\{L,R\}$, let
$\Lambda^{X}_{2,O}\triangleq\bigcup_{i\in\{1,3,5\}}\Lambda^{X}_{2,i}$
and
$\Lambda^{X}_{2,E}\triangleq\bigcup_{i\in\{2,4,6\}}\Lambda^{X}_{2,i}.$
Correspondingly, for $X\in\{L,R\}$ and $Y\in\{O,E\}$, let
$\Omega^{X}_{2,Y}\triangleq\bigcup_{(\bm{z},\bm{z}^{\prime})\in\Lambda^{X}_{2,Y}}B^{\text{S}}_{1}(\bm{z},\bm{z}^{\prime})$
and
$\Omega_{2,Y}\triangleq\Omega^{L}_{2,Y}\cup\Omega^{R}_{2,Y}.$
Lemma 3
Suppose $d=2$. The following hold.
  • 1)
    $\left|\Lambda^{L}_{2,O}\cup\Lambda^{R}_{2,O}\right|\leq n-2$.
  • 2)
    For each $X\in\{L,R\}$, if $|T^{X}\cap[i_{1}+1,i_{2}]|\neq 0$, then $|\Lambda^{X}_{2,E}|\leq 6$.
  • 3)
    For each $X\in\{L,R\}$, if $|T^{X}\cap[i_{1}+1,i_{2}]|=0$, then $|\Omega^{X}_{2,E}\backslash\Omega^{X}_{0}|\leq 6$.
  • 4)
    If $|T^{L}\cap[i_{1}+1,i_{2}]|=|T^{R}\cap[i_{1}+1,i_{2}]|=0$, then we have $|\Omega_{0}|=2(1+(q-1)(n-1))-q=2(q-1)n-3q+2$ and $|\Omega_{2}\backslash\Omega_{0}|\leq 2n-6-\delta_{q,2}$.
proof 13
1) Since $d=2$, then from Claims 2.1, 2.3, 2.5 and Claims 2.1, 2.3, 2.5, we can obtain $\Lambda^{L}_{2,O}=\Lambda^{R}_{2,O}=\big{\{}(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j}):j\in[n]\backslash\{i_{1},i_{2}\}\big{\}}.$ Hence,
$|\Lambda^{L}_{2,O}\cup\Lambda^{R}_{2,O}|=r_{1}+r_{2}+r_{3}\leq n-2$
where $r_{1}$ is the number of runs of $x_{[1,i_{1}-1]}$, $r_{2}$ is the number of runs of $x_{[i_{2}+1,n]}$ and $r_{3}$ is the number of runs of $x_{[i_{1}+1,i_{2}-1]}$.
2) Note that for $d=2$, we have $[i_{1}+1,i_{d-1}]=[i_{2}+1,i_{d}]=\emptyset$. Then this statement can be obtained directly from Claims 2.2, 2.4, 2.6 and Claims 2.2, 2.4, 2.6.
3) As the proof for $X=R$ and for $X=L$ are similar, we only prove the result for $X=L$. Denote $(\bm{z}_{1},\bm{z}_{1}^{\prime})\triangleq(x_{[n]\backslash k_{1}-1},x^{\prime}_{[n]\backslash k^{\prime}_{1}})$, $(\bm{z}_{2},\bm{z}_{2}^{\prime})\triangleq(x_{[n]\backslash k_{1}-1},x^{\prime}_{[n]\backslash i_{1}})$ and $(\bm{z}_{3},\bm{z}_{3}^{\prime})\triangleq(x_{[n]\backslash i_{2}},x^{\prime}_{[n]\backslash k^{\prime}_{1}})$. Note that $d=2$. By checking Claims 2.2, 2.4 and 2.6, we can find that for each $(\bm{z},\bm{z}^{\prime})\in\Lambda^{L}_{2,E}\backslash\{(\bm{z}_{i},\bm{z}_{i}^{\prime}):i\in\{1,2,3\}\}$, either $\bm{z}=x_{[n]\backslash i_{1}}$ or $\bm{z}^{\prime}=x^{\prime}_{[n]\backslash i_{2}}$, so by Remark 6, we have $B^{\text{S}}_{1}(\bm{z},\bm{z}^{\prime})\subseteq\Omega^{L}_{0}$. Therefore, we can obtain $\Omega^{L}_{2,E}\backslash\Omega^{L}_{0}\subseteq\bigcup_{i=1}^{3}B^{\text{S}}_{1}(\bm{z}_{i},\bm{z}_{i}^{\prime})\subseteq\Lambda^{L}_{2}$, and so $|\Omega^{L}_{2,E}\backslash\Omega^{L}_{0}|\leq\sum_{i=1}^{3}|B^{\text{S}}_{1}(\bm{z}_{i},\bm{z}_{i}^{\prime})|\leq 6$.
4) Since $d=2$ and $|T^{L}\cap[i_{1}+1,i_{2}]|=|T^{R}\cap[i_{1}+1,i_{2}]|=0$, we have $i_{2}=i_{1}+1$. ((((In fact, if $i_{2}>i_{1}+1$, then we can obtain $x_{i_{1}}=x^{\prime}_{i_{1}+1}=x_{i_{1}+1}=x^{\prime}_{i_{1}}$, which contradicts to the definition of $S.)$ So, $\bm{x}$ and $\bm{x}^{\prime}$ are of the form
$\bm{x}$ $=\bm{u}ab\bm{v}$
$~{}\bm{x}^{\prime}$ $=\bm{u}ba\bm{v}$
where $a=x_{i_{1}}\neq b=x_{i_{2}}$, $\bm{u}\in\Sigma_{q}^{i_{1}-1}$ and $\bm{v}\in\Sigma_{q}^{n-i_{2}}$.
By Claims 0 and 0, we have $\Lambda_{0}=\big{\{}(x_{[n]\backslash i_{1}},x^{\prime}_{[n]\backslash i_{2}})$, $(x_{[n]\backslash i_{2}},x^{\prime}_{[n]\backslash i_{1}})\big{\}}$, so we can obtain $\Omega_{0}=B^{\text{S}}_{1}(x_{[n]\backslash i_{1}})\cup B^{\text{S}}_{1}(x_{[n]\backslash i_{2}})$. Note that $d_{\text{H}}(x_{[n]\backslash i_{1}},x_{[n]\backslash i_{2}})=1$. Then $|\Omega_{0}|=|B^{\text{S}}_{1}(x_{[n]\backslash i_{1}})|+|B^{\text{S}}_{1}(x_{[n]\backslash i_{2}})-|B^{\text{S}}_{1}(x_{[n]\backslash i_{1}})\cap B^{\text{S}}_{1}(x_{[n]\backslash i_{2}})|=2(1+(q-1)(n-1))-q=2(q-1)n-3q+2$.
In the following, we prove that $|\Omega_{2}\backslash\Omega_{0}|\leq 2n-6-\delta_{q,2}$.
By Claims 2.2, 2.4, 2.6 and Claims 2.2, 2.4, 2.6, we find:
  • $\Lambda^{L}_{2,2}=\big{\{}(x_{[n]\backslash k_{2}-1},x^{\prime}_{[n]\backslash i_{2}}),(x_{[n]\backslash k_{1}-1},x^{\prime}_{[n]\backslash k^{\prime}_{1}})$, $(x_{[n]\backslash i_{1}},x^{\prime}_{[n]\backslash k^{\prime}_{2}})\big{\}}$ and $\Lambda^{R}_{2,2}=\big{\{}(x_{[n]\backslash i_{2}},x^{\prime}_{[n]\backslash m_{2}-1})$, $(x_{[n]\backslash m^{\prime}_{1}},x^{\prime}_{[n]\backslash m_{1}-1}),(x_{[n]\backslash m^{\prime}_{2}},x^{\prime}_{[n]\backslash i_{1}})\big{\}}$.
  • $\Lambda^{L}_{2,4}=\big{\{}(x_{[n]\backslash k_{1}-1},x^{\prime}_{[n]\backslash i_{1}})\big{\}}$ and $\Lambda^{R}_{2,4}=\big{\{}(x_{[n]\backslash i_{1}},x^{\prime}_{[n]\backslash m_{1}-1})\big{\}}$. Note that $d=2$ and $i_{1}=i_{2}-1$, so $T^{L}\cap[i_{d-1}+1,i_{d}-1]=\emptyset$ and $j_{1}^{\prime}$ does not exists.
  • $\Lambda^{L}_{2,6}=\big{\{}(x_{[n]\backslash i_{2}},x^{\prime}_{[n]\backslash k^{\prime}_{1}})\big{\}}$ and $\Lambda^{R}_{2,6}=\big{\{}(x_{[n]\backslash m^{\prime}_{1}},x^{\prime}_{[n]\backslash i_{2}})\big{\}}$. Note that $i_{1}=i_{2}-1$, so $T^{L}\cap[i_{1}+2,i_{2}]=\emptyset$ and $j_{1}$ does not exists.
Here $k_{1}$, $k^{\prime}_{1}$, $m_{1}$, $m_{1}^{\prime}$ are defined by (7)--(10), and $k_{2}$, $k^{\prime}_{2}$, $m_{2}$, $m_{2}^{\prime}$ are defined by (14)--(17), respectively.
For each $(\bm{z},\bm{z}^{\prime})\in\Lambda^{L}_{2,E}\cup\Lambda^{R}_{2,E}$ except for $(\bm{u}_{1},\bm{u}_{1}^{\prime})\triangleq(x_{[n]\backslash k_{1}-1},x^{\prime}_{[n]\backslash k^{\prime}_{1}})$ and $(\bm{u}_{2},\bm{u}_{2}^{\prime})\triangleq(x_{[n]\backslash m^{\prime}_{1}},x^{\prime}_{[n]\backslash m_{1}-1})$, we find that either $\bm{z}\in\{x_{[n]\backslash i_{1}},x_{[n]\backslash i_{2}},x^{\prime}_{[n]\backslash i_{1}},x^{\prime}_{[n]\backslash i_{2}}\}$ or $\bm{z}^{\prime}\in\{x_{[n]\backslash i_{1}},x_{[n]\backslash i_{2}},x^{\prime}_{[n]\backslash i_{1}},x^{\prime}_{[n]\backslash i_{2}}\}$, so by Remark 6, we have $B^{\text{S}}_{1}(\bm{z},\bm{z}^{\prime})\subseteq\Omega_{0}$, which implies that
$(\Omega^{L}_{2,E}\cup\Omega^{R}_{2,E})\backslash\Omega_{0}\subseteq\bigcup_{i=1}^{2}B^{\text{S}}_{1}(\bm{u}_{i},\bm{u}^{\prime}_{i}).$
For $(\bm{u}_{1},\bm{u}_{1}^{\prime})=(x_{[n]\backslash k_{1}-1},x^{\prime}_{[n]\backslash k^{\prime}_{1}})$, we have
$B^{\text{S}}_{1}(x_{[n]\backslash k_{1}-1},x^{\prime}_{[n]\backslash k^{\prime}_{1}})=\left\{\phi^{k_{1}-1}_{k_{1};x^{\prime}_{k_{1}-1}}(\bm{x}),\phi^{k_{1}-1}_{k^{\prime}_{1};x^{\prime}_{k^{\prime}_{1}-1}}(\bm{x})\right\}.$
By the definition of $k_{1}$, we can find $x_{k_{1}-1}=x^{\prime}_{k_{1}-1}~{}($because $k_{1}-1<i_{1})$, and so
$\phi^{k_{1}-1}_{k_{1};x^{\prime}_{k_{1}-1}}(\bm{x})=x_{[n]\backslash k_{1}}.$
We can also find $x_{i}=x^{\prime}_{i}=x_{i+1}$ for any $i\in[k_{1},i_{1}-1]$, so
$\phi^{k_{1}-1}_{k_{1};x^{\prime}_{k_{1}-1}}(\bm{x})=x_{[n]\backslash k_{1}}=x_{[n]\backslash i_{1}}\in\Omega_{0}.$
Similarly, for $(\bm{u}_{2},\bm{u}_{2}^{\prime})=(x_{[n]\backslash m^{\prime}_{1}},x^{\prime}_{[n]\backslash m_{1}-1})$, we have
$B^{\text{S}}_{1}(\bm{u}_{2},\bm{u}_{2}^{\prime})=\left\{\phi^{m^{\prime}_{1}}_{m^{\prime}_{1}-1;x^{\prime}_{m^{\prime}_{1}}}(\bm{x}),\phi^{m^{\prime}_{1}}_{m_{1}-1;x^{\prime}_{m_{1}}}(\bm{x})\right\},$
and by the definition of $m_{1}^{\prime}$, we can find
$\phi^{m^{\prime}_{1}}_{m^{\prime}_{1}-1;x^{\prime}_{m^{\prime}_{1}}}(\bm{x})=x_{[n]\backslash m^{\prime}_{1}-1}=x_{[n]\backslash i_{2}}\in\Omega_{0}.$
Thus, we have
$\Omega_{2,E}\backslash\Omega_{0}\subseteq\Big{\{}\phi^{k_{1}-1}_{k^{\prime}_{1};x^{\prime}_{k^{\prime}_{1}-1}}(\bm{x}),\phi^{m^{\prime}_{1}}_{m_{1}-1;x^{\prime}_{m_{1}}}(\bm{x})\Big{\}}.$
Consider $\Lambda_{2,O}\triangleq\bigcup_{i\in\{1,3,5\}}(\Lambda^{L}_{2,i}\cup\Lambda^{R}_{2,i})$. By Claims 2.1, 2.3, 2.5 and Claims 2.1, 2.3, 2.5, we can obtain:
  • $\Lambda^{L}_{2,1}=\Lambda^{R}_{2,1}=\big{\{}(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j}):j\in[i_{2}+1,n]\big{\}}$.
  • $\Lambda^{L}_{2,3}=\Lambda^{R}_{2,3}=\big{\{}(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j}):j\in[1,i_{1}-1]\big{\}}$.
  • $\Lambda^{L}_{2,5}=\Lambda^{R}_{2,5}=\emptyset$.
So, we can obtain
$\Omega_{2,O}$ =$=\big{\{}B^{\text{S}}_{1}(x_{[n]\backslash j})\cup B^{\text{S}}_{1}(x^{\prime}_{[n]\backslash j}):j\in[n]\backslash\{i_{1},i_{2}\}\big{\}}$
$=\bigcup_{j\in[n]\backslash\{i_{1},i_{2}\}}\big{\{}\phi^{j}_{i_{1};x^{\prime}_{i_{1}}}(\bm{x}),\phi^{j}_{i_{2};x^{\prime}_{i_{2}}}(\bm{x})\big{\}},$
where the second equation holds because for each $j\in[n]\backslash\{i_{1},i_{2}\}$, $x_{[n]\backslash j}$ and $x^{\prime}_{[n]\backslash j}$ are of the form
$x_{[n]\backslash j}$ $=\bm{u}^{\prime}ab\bm{v}^{\prime}$
$~{}x^{\prime}_{[n]\backslash j}$ $=\bm{u}^{\prime}ba\bm{v}^{\prime}$
where $\bm{u}^{\prime}\bm{v}^{\prime}$ is obtained from $\bm{u}\bm{v}$ by a single deletion. For $j=i_{1}-1$, we find
$\phi^{i_{1}-1}_{i_{1};x^{\prime}_{i_{1}}}(\bm{x})=\phi^{i_{1}}_{i_{1}-1;x^{\prime}_{i_{1}}}(\bm{x})\in\Omega_{0}.$
Moreover, we find
$\phi^{i_{1}-1}_{i_{2};x^{\prime}_{i_{2}}}(\bm{x})$ $=\bm{u}^{\prime}aa\bm{v}$
$~{}x_{[n]\backslash i_{2}}$ $=\bm{u}^{\prime}ca\bm{v}$
where $\bm{u}^{\prime}=x_{[1,i_{1}-2]}$ and $c=x_{i_{1}-1}$. Hence, we have $\phi^{i_{1}-1}_{i_{2};x^{\prime}_{i_{2}}}(\bm{x})\in B^{\text{S}}_{1}(x_{[n]\backslash i_{2}})\subseteq\Omega_{0}$, which implies that $\{\phi^{i_{1}-1}_{i_{1};x^{\prime}_{i_{1}}}(\bm{x}),\phi^{i_{1}-1}_{i_{2};x^{\prime}_{i_{2}}}(\bm{x})\}\subseteq\Omega_{0}$. Similarly, we can prove that $\{\phi^{i_{2}+1}_{i_{1};x^{\prime}_{i_{1}}}(\bm{x}),\phi^{i_{2}+1}_{i_{2};x^{\prime}_{i_{2}}}(\bm{x})\}\subseteq\Omega_{0}$. Thus,
$\Omega_{2,O}=\bigcup_{j\in[n]\backslash[i_{1}-1,i_{2}+1]}\{\phi^{j}_{i_{1};x^{\prime}_{i_{1}}}(\bm{x}),\phi^{j}_{i_{2};x^{\prime}_{i_{2}}}(\bm{x})\}$
and so
$|\Omega_{2,E}\backslash\Omega_{0}|+|\Omega_{2,O}\backslash\Omega_{0}|\leq 2+2(n-4)=2n-6.$
Note that if $i_{1}=1$ or $i_{2}=n$, then $k_{1}$ and $m_{1}^{\prime}$ do not exist and so $\phi^{k_{1}-1}_{k^{\prime}_{1};x^{\prime}_{k^{\prime}_{1}-1}}(\bm{x})$ and $\phi^{m^{\prime}_{1}}_{m_{1}-1;x^{\prime}_{m_{1}}}(\bm{x})$ do not exist. Thus, $|\Omega_{2,E}\backslash\Omega_{0}|=0$ and $|\Omega_{2,O}\backslash\Omega_{0}|\leq 2(n-3)$, and so still
$|\Omega_{2,E}\backslash\Omega_{0}|+|\Omega_{2,O}\backslash\Omega_{0}|\leq 2n-6.$
If $p=2$, consider $j=i_{1}-2$. We find that
$\phi^{i_{1}-2}_{i_{1};x^{\prime}_{i_{1}}}(\bm{x})$ $=\bm{u}^{\prime\prime}cbb\bm{v}$
$~{}x_{[n]\backslash i_{1}}$ $=\bm{u}^{\prime\prime}ecb\bm{v}$
and
$\phi^{i_{1}-2}_{i_{2};x^{\prime}_{i_{2}}}(\bm{x})$ $=\bm{u}^{\prime\prime}caa\bm{v}$
$~{}x_{[n]\backslash i_{2}}$ $=\bm{u}^{\prime\prime}eca\bm{v}$
where $\bm{u}^{\prime\prime}=x_{[1,i_{1}-3]}$, $c=x_{i_{1}-1}$ and $e=x_{i_{1}-2}$. Since $p=2$, then either $c=a$ or $c=b$. Hence, either $d_{\text{H}}(\phi^{i_{1}-2}_{i_{1};x^{\prime}_{i_{1}}}(\bm{x}),x_{[n]\backslash i_{1}})=1$ or $d_{\text{H}}(\phi^{i_{1}-2}_{i_{2};x^{\prime}_{i_{2}}}(\bm{x}),x_{[n]\backslash i_{2}})=1$, and so either $\phi^{i_{1}-2}_{i_{1};x^{\prime}_{i_{1}}}(\bm{x})\in B^{\text{S}}_{1}(x_{[n]\backslash i_{1}})\subseteq\Omega_{0}$ or $\phi^{i_{1}-2}_{i_{2};x^{\prime}_{i_{2}}}(\bm{x})\in B^{\text{S}}_{1}(x_{[n]\backslash i_{2}})\subseteq\Omega_{0}$, which implies that |$|\{\phi^{i_{1}-2}_{i_{1};x^{\prime}_{i_{1}}}(\bm{x}),\phi^{i_{1}-2}_{i_{2};x^{\prime}_{i_{2}}}(\bm{x})\}\backslash\Omega_{0}|=1$. Similarly, we can prove that |$|\{\phi^{i_{2}+2}_{i_{1};x^{\prime}_{i_{1}}}(\bm{x}),\phi^{i_{2}+2}_{i_{2};x^{\prime}_{i_{2}}}(\bm{x})\}\backslash\Omega_{0}|=1$. So,
$|\Omega_{2,E}\backslash\Omega_{0}|+|\Omega_{2,O}\backslash\Omega_{0}|\leq 2n-6-2=2n-8.$
On the other hand, if $i_{1}=2~{}($or $i_{2}=n-1)$, by $q=2$ and $x_{1}=x^{\prime}_{1}~{}($resp. $x_{n}=x_{n}^{\prime})$, we can find $x_{1}=x^{\prime}_{2}$ or $x^{\prime}_{1}=x_{2}~{}($resp. $x_{n}=x^{\prime}_{n-1}$ or $x_{n-1}=x^{\prime}_{n})$, and so $k_{1}$ or $m_{1}$ does not exists ((((resp. $k^{\prime}_{1}$ or $m^{\prime}_{1}$ does not exists)))), which implies that still $|\Omega_{2,E}\backslash\Omega_{0}|+|\Omega_{2,O}\backslash\Omega_{0}|\leq 2n-8~{}($because $|\Omega_{2,E}\backslash\Omega_{0}|=1$ and $|\Omega_{2,O}\backslash\Omega_{0}|\leq 2n-9)$. If $i_{1}=1$ or $i_{2}=n$, then we have $|\Omega_{2,E}\backslash\Omega_{0}|+|\Omega_{2,O}\backslash\Omega_{0}|\leq 2n-7~{}($because $|\Omega_{2,E}\backslash\Omega_{0}|=0$ and $|\Omega_{2,O}\backslash\Omega_{0}|\leq 2n-7)$.
By the above discussions, we can obtain $|\Omega_{2}\backslash\Omega_{0}|\leq 2n-6-\delta_{q,2}$, where $\delta_{q,2}=1$ if $q=2$, and $\delta_{q,2}=0$ otherwise.
Lemma 4
Suppose $d\geq 3$. The following hold.
  • 1)
    For each $X\in\{L,R\}$, if $|T^{X}\cap[i_{1}+1,i_{d}]|=0$, then $|\Omega^{X}_{2}\backslash\Omega^{X}_{0}|\leq 8$.
  • 2)
    For each $X\in\{L,R\}$, if $|T^{X}\cap[i_{1}+1,i_{d}]|\neq 0$, then $|\Lambda^{X}_{2}|\leq 8$.
proof 14
1) For $X=L$, the result can be obtained from Claims 2.1--2.6 and Remark 6; for $X=R$, the result can be obtained from Claims ${}^{\prime}-$2.6 and Remark 6.
2) The result can be obtained directly from Claims 2.1--2.6 ((((for $X=L)$ and Claims ${}^{\prime}-$${}^{\prime}~{}($for $X=R)$.
Now, we prove the upper bound in Theorem 1, that is, if $d=d_{\text{H}}(\bm{x},\bm{x}^{\prime})\geq 2$ and $n\geq\max\{\frac{q+23}{2},\frac{5q+19}{q-1}\}$, then
$|B^{\text{D,S}}_{1,1}(\bm{x},\bm{x}^{\prime})|\leq 2qn-3q-2-\delta_{q,2}.$
Suppose $d\geq 3$.
  • If $|T^{L}\cap[i_{1}+1,i_{d}]|=|T^{R}\cap[i_{1}+1,i_{d}]|=0$, then by Claim 0, Claim 0, 1) of Lemma 2, and 1) of Lemma 4, we have $|B^{\text{D,S}}_{1,1}(\bm{x},\bm{x}^{\prime})|\leq 2(1+(q-1)(n-1))+2(8)=2(q-1)n-2q+20\leq 2qn-3q-2-\delta_{q,2}$, where the last inequality comes from the assumption that $n\geq\max\{\frac{q+23}{2},\frac{5q+19}{q-1}\}$.
  • If $|T^{L}\cap[i_{1}+1,i_{d}]|=0$ and $|T^{R}\cap[i_{1}+1,i_{d}]|\neq 0$, then by Claim 0, 1) of Lemma 2, 1) of Lemma 4; 3) of Lemma 2, and 2) of Lemma 4, we have $|B^{\text{D,S}}_{1,1}(\bm{x},\bm{x}^{\prime})|\leq(1+(q-1)(n-1))+8+q(2)+2(8)=(q-1)n+q+26\leq 2qn-3q-2-\delta_{q,2}$.
  • If $|T^{L}\cap[i_{1}+1,i_{d}]|\neq 0$ and $|T^{R}\cap[i_{1}+1,i_{d}]|=0$, then by Claim 0, 1) of Lemma 2, 1) of Lemma 4, 3) of Lemma 2, and 2) of Lemma 4, we have $|B^{\text{D,S}}_{1,1}(\bm{x},\bm{x}^{\prime})|\leq(1+(q-1)(n-1))+8+q(2)+2(8)=(q-1)n+q+26\leq 2qn-3q-2-\delta_{q,2}$.
  • If $|T^{L}\cap[i_{1}+1,i_{d}]|\neq 0$ and $|T^{R}\cap[i_{1}+1,i_{d}]|\neq 0$, then by Claim 0, Claim 0, 3) of Lemma 2, and 2) of Lemma 4, we have $|B^{\text{D,S}}_{1,1}(\bm{x},\bm{x}^{\prime})|\leq 2(q(2)+2(8))=4q+32\leq 2qn-3q-2-\delta_{q,2}$.
Suppose $d=2$.
  • If $|T^{L}\cap[i_{1}+1,i_{2}]|=|T^{R}\cap[i_{1}+1,i_{2}]|=0$, then by Claim 0, Claim 0, 1) of Lemma 2 and 4) of Lemma 3, we have $|B^{\text{D,S}}_{1,1}(\bm{x},\bm{x}^{\prime})|\leq 2(1+(q-1)(n-1))-q+2n-6-\delta_{q,2}=2qn-3q-2-\delta_{q,2}$.
  • If $|T^{L}\cap[i_{1}+1,i_{2}]|=0$ and $|T^{R}\cap[i_{1}+1,i_{2}]|\neq 0$, then by Claim 0, 1) of Lemma 2, 3) of Lemma 3; 2) of Lemma 2, and 1)--2) of Lemma 3, we have $|B^{\text{D,S}}_{1,1}(\bm{x},\bm{x}^{\prime})|\leq(1+(q-1)(n-1))+6+q(3)+2(n-2+6)=(q+1)n+2q+16\leq 2qn-3q-2-\delta_{q,2}$, where the last inequality comes from the assumption that $n\geq\max\{\frac{q+23}{2},\frac{5q+19}{q-1}\}$.
  • If $|T^{L}\cap[i_{1}+1,i_{2}]|\neq 0$ and $|T^{R}\cap[i_{1}+1,i_{2}]|=0$, then by Claim 0, 1) of Lemma 2, 3) of Lemma 3; 2) of Lemma 2, and 1)--2) of Lemma 3, we have $|B^{\text{D,S}}_{1,1}(\bm{x},\bm{x}^{\prime})|\leq(1+(q-1)(n-1))+6+q(3)+2(n-2+6)=(q+1)n+2q+16\leq 2qn-3q-2-\delta_{q,2}$.
  • If $|T^{L}\cap[i_{1}+1,i_{D}]|\neq 0$ and $|T^{R}\cap[i_{1}+1,i_{0}]|\neq 0$, then by 2) of Lemma 2, and 1)--2) of Lemma 3, we have $|B^{\text{D,S}}_{1,1}(\bm{x},\bm{x}^{\prime})|\leq 2q(3)+2(n-2+6+6)=2n+6q+20\leq 2qn-3q-2-\delta_{q,2}$.
Thus, we can obtain $|B^{\text{D,S}}_{1,1}(\bm{x},\bm{x}^{\prime})|\leq 2qn-3q-2-\delta_{q,2}$.
To prove the tightness of this bound, we give two examples of $\bm{x},\bm{x}^{\prime}\in\Sigma_{q}^{n}$ such that $d_{\text{H}}(\bm{x},\bm{x}^{\prime})=2$ and $|B^{\text{D,S}}_{1,1}(\bm{x},\bm{x}^{\prime})|=2qn-3q-2-\delta_{q,2}$.
Example 2
Let $q\geq 3$ and $n\geq 5$. Let $\bm{x},\bm{x}^{\prime}\in\Sigma_{q}^{n}$ such that
$\bm{x}$ $=01201A_{n-5}(01)$
$\bm{x}^{\prime}$ $=10201A_{n-5}(01).$
We have $S=\{i_{1},i_{2}\}=\{1,2\}$, where $S$ is defined according to (3). It is not hard to verify that
  • $\Omega_{0}=B^{\text{S}}_{1}(x_{[n]\backslash 1})\cup B^{\text{S}}_{1}(x_{[n]\backslash 2})$ and $d_{\text{H}}(x_{[n]\backslash 1},x^{\prime}_{[n]\backslash 2})=1$, where $x_{[n]\backslash 1}=1201A_{n-5}(01)$ and $x_{[n]\backslash 2}=0201A_{n-5}(01)$, so $|\Omega_{0}|=2(1+(q-1)(n-1))-q=2(q-1)n-3q+4$.
  • Let $\Omega_{2}^{\prime}\triangleq\bigcup_{j=4}^{n}\big{\{}\phi^{j}_{1;1}(\bm{x}),\phi^{j}_{2;0}(\bm{x})\big{\}}$. Then $\Omega_{2}^{\prime}\subseteq\Omega_{2,O}$ and $\big{|}\Omega_{2}^{\prime}|=2(n-3)$. Moreover, for each $\bm{z}\in\Omega_{2}^{\prime}$ and each $\bm{z}^{\prime}\in\{x_{[n]\backslash 1},x_{[n]\backslash 2}\}$, we have $d_{\text{H}}(\bm{z},\bm{z}^{\prime})\geq 2$, and so $\Omega_{2}^{\prime}\cap\Omega_{0}=\emptyset$, which implies that $\Omega_{2}^{\prime}\subseteq\Omega_{2,O}\backslash\Omega_{0}$.
Thus, we can obtain
$|B^{\text{D,S}}_{1,1}(\bm{x},\bm{x}^{\prime})|\geq|\Omega_{0}|+|\Omega_{2}^{\prime}|=2qn-3q-2.$
Example 3
Let $q=2$ and $n\geq 4$. Let $\bm{x},\bm{x}^{\prime}\in\Sigma_{q}^{n}$ such that
$\bm{x}$ $=0101A_{n-4}(01)$
$\bm{x}^{\prime}$ $=1001A_{n-4}(01).$
It is not hard to verify that
  • $\Omega_{0}=B^{\text{S}}_{1}(x_{[n]\backslash 1})\cup B^{\text{S}}_{1}(x_{[n]\backslash 2})$ and $d_{\text{H}}(x_{[n]\backslash 1},x^{\prime}_{[n]\backslash 2})=1$, so $|\Omega_{0}|=2(1+(q-1)(n-1))-q=2(q-1)n-3q+4$.
  • Let $\Omega_{2}^{\prime}\triangleq\bigcup_{j=4}^{n}\big{\{}\phi^{j}_{1;1}(\bm{x}),\phi^{j}_{2;0}(\bm{x})\big{\}}\backslash\{\phi^{4}_{2;0}(\bm{x})\}$. Then $\Omega_{2}^{\prime}\subseteq\Omega_{2}$ and $\big{|}\Omega_{2}^{\prime}\big{|}=2(n-3)-1$. Moreover, for each $\bm{z}\in\Omega_{2}^{\prime}$ and each $\bm{z}^{\prime}\in\{x_{[n]\backslash 1},x_{[n]\backslash 2}\}$, we have $d_{\text{H}}(\bm{z},\bm{z}^{\prime})\geq 2$, and so $\Omega_{2}^{\prime}\subseteq\Omega_{2,O}\backslash\Omega_{0}$. Note that in this example, we can see that $\phi^{4}_{2;0}(\bm{x})=000A_{n-4}(01)$ and $x_{[n]\backslash 2}=001A_{n-4}(01)$, so $\phi^{4}_{2;0}(\bm{x})\in B^{\text{S}}_{1}(x_{[n]\backslash 2})\subseteq\Omega_{0}$.
Hence, we have
$|B^{\text{D,S}}_{1,1}(\bm{x},\bm{x}^{\prime})|\geq|\Omega_{0}|+|\Omega_{2}^{\prime}|=2qn-3q-3=4n-9.$
By the above discussions, we prove Theorem 1.
Remark 7
By the definition of $T^{L}$ and $T^{R}$, it is easy to see that if the Levenshtein distance $d_{\text{L}}(\bm{x},\bm{x}^{\prime})\geq 2$, then we must have $|T^{L}\cap[i_{1}+1,i_{d}]|\neq 0$ and $|T^{R}\cap[i_{1}+1,i_{d}]|\neq 0$. Therefore, if $d_{\text{H}}(\bm{x},\bm{x}^{\prime})\geq 3$ and $d_{\text{L}}(\bm{x},\bm{x}^{\prime})\geq 2$, then by the proof of Theorem 1, we can obtain $|B^{\text{D,S}}_{1,1}(\bm{x},\bm{x}^{\prime})|\leq 4q+32$, which depends only on $q$. However, this bound is not tight. To obtain a tight bound independent of $n$ for this case, more careful discussions are needed. This problem will be investigated in our future work.
V Conclusions and future work
We proved a tight upper bound on the intersection size of error balls of single-deletion single-substitution channel for any $q$-ary sequences $\bm{x},\bm{x}^{\prime}$ of length $n$ and with Hamming distance $d_{\text{H}}(\bm{x},\bm{x}^{\prime})\geq 2$. This upper bound is the minimum number of channel outputs ((((reads)))) required to reconstruct a sequence in a code with minimum Hamming distance $2$.
The bound obtained in this work depends on the sequence length $n$. If we consider any $\bm{x},\bm{x}^{\prime}\in\Sigma_{q}^{n}$ with Hamming distance $d_{\text{H}}(\bm{x},\bm{x}^{\prime})\geq 3$ and Levenshtein distance $d_{\text{L}}(\bm{x},\bm{x}^{\prime})\geq 2$, then as pointed out in Remark 7, we can obtain an upper bound of $|B^{\text{D,S}}_{1,1}(\bm{x},\bm{x}^{\prime})|$ depending only on $q$. For binary code, this requirement can be satisfied by introducing a redundancy of only $\log n$ bits. The problem of constructing reconstruction codes with constant number of reads (i.e., the number of reads is independent of $n$ and depend only on $q$) for single-deletion single-substitution channel is left in our future work.